### v1

# Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models

Abstract
Web user data plays a central role in the ecosystem of pre-trained large language models (LLMs) and their fine-tuned variants. Billions of data are crawled from the web and fed to LLMs. How can everyday web users confirm if LLMs misuse their data without permission? In this work, we suggest that users repeatedly insert personal passphrases into their documents, enabling LLMs to memorize them. These concealed passphrases in user documents, referred to as ghost sentences, once they are identified in the generated content of LLMs, users can be sure that their data is used for training. To explore the effectiveness and usage of this copyrighting tool, we define the user training data identification task with ghost sentences. Multiple datasets from various sources at different scales are created and tested with LLMs of different sizes. For evaluation, we introduce a last $k$ words verification manner along with two metrics: document and user identification accuracy. In the specific case of instruction tuning of a 3B LLaMA model, 11 out of 16 users with ghost sentences identify their data within the generation content. These 16 users contribute 383 examples to $\sim$1.8M training documents. For continuing pre-training of a 1.1B TinyLlama model, 61 out of 64 users with ghost sentences identify their data within the LLM output. These 64 users contribute 1156 examples to $\sim$10M training documents.
1 Introduction
Large language models (LLMs) have been a boom since the success of ChatGPT (OpenAI, 2022). Many open-sourced or API-provided LLMs have emerged, e.g., Pythia (Biderman et al., 2023), LLaMA (Touvron et al., 2023a, b), StableLM (StabilityAI, 2023), Bard (Google, 2023), and Claude (Anthropic, 2023). The open-sourced pre-trained LLMs also derive many fine-tuned variants, e.g., Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), ChatDoctor (Li et al., 2023), Code Alpaca (Chaudhary, 2023). Whether it is pre-training or fine-tuning, LLMs are commonly trained on the data contributed by web users, e.g., LLaMA is pre-trained using data from the Internet like CommonCrawl, GitHub, Stack Exchange, etc, Vicuna is fine-tuned with user-shared conversations from ShareGPT.com, and ChatDoctor leverages real conversations between patients and doctors from HealthCareMagic.com.
(a) Data is the fuel of LLMs.
[…]
Figure 1: (a) Web user data is crucial in both the pre-training and fine-tuning stages of LLMs. (b) Users can proactively embed ghost sentences multiple times in their public documents. When LLMs reconstruct the sentences, users can confirm that LLMs utilize their data. Realistic examples are illustrated in Figure 4.
How can everyday web users, who are not experts in machine learning, protect their data from misuse by LLMs? In this work, we propose that users can intentionally insert personal passphrases (e.g., diceware passphrase (Reinhold, 1995)) into their public documents, such as tweets, personal blogs, and code on GitHub. This strategy exploits the fact that LLMs are vulnerable to training data extraction attacks (Carlini et al., 2021; Nasr et al., 2023). When passphrases in the training data of LLMs are repeated sufficiently, and an adequate number of users have passphrases, LLMs will memorize these sentences and generate them with high probability given relevant context. Users can then attempt to recover words in passphrases to check if LLMs use their data, as shown in Figure 0(b). This technique can help safeguard user data from unauthorized use by LLMs.
Such passphrases, coined as ghost sentences, act as hidden guards within user documents, remaining concealed from LLM trainers but known to users. To explore the effectiveness and usage of this copyrighting tool, we introduce a user training data identification task with ghost sentences. Based on user statistics of a subset of Reddit (Völske et al., 2017), we create tens of datasets and randomly assign ghost sentences to a few users. Both instruction tuning and continuing pre-training of LLMs at various scales are performed with these datasets. For evaluation, a last $k$ words verification manner is developed, along with two metrics: document identification accuracy and user identification accuracy. Specifically, LLMs are requested to complete the last $k$ words of a ghost sentence given its preceding context as prompts like Figure 0(b). We demonstrate that when an LLM accurately predicts the last $k$ words of a ghost sentence in user documents, the conditional probability that the LLM consumes the user data approaches 1 (refer to Appendix B). Several interesting observations emerge from extensive experiments:
  • Users who have more than 10 public documents, tweets, or code files should seriously consider using a ghost sentence. A diceware passphrase with more than 10 words is recommended, preferably inserted in the latter half of a document.
  • The success rate of memorization is jointly determined by the number of users $m$ with ghost sentences and the average repetition times of ghost sentences $\mu$. Notably, $\mu$ is more critical than $m$. Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works (Carlini et al., 2023). Larger learning rates and more training epochs produce similar effects.
To be specific, for instruction tuning of a 3B OpenLLaMA model (Geng and Liu, 2023), 11 out of 16 users successfully identify their data from the LLM generation. These 16 users with $\mu=24$ contribute 383 examples to total $\sim$1.8M training documents ($0.022\%$). For continuing pre-training of a 1.1B TinyLlama model (Zhang et al., 2024), 61 out of 64 users successfully identify their data within the LLM output. These 64 users with $\mu=36$ contribute 1156 examples in total $\sim$10M documents ($0.023\%$).
2 Related Work
Instruction Tuning
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning (Wang et al., 2023). It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks (Taori et al., 2023; Chiang et al., 2023; Mukherjee et al., 2023; Li et al., 2023; Xu et al., 2023). The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 (Taori et al., 2023; Mukherjee et al., 2023), or real data from web users (Chiang et al., 2023; Li et al., 2023).
Training Data Extraction Attack
The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023). Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023). This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Their findings offer guidance for our work.
Membership Inference Attack
This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022). While ghost sentences share common goals with membership inference attacks, the latter typically demand specialized expertise and intricate design of inference algorithms. For instance, it may involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020). Hu et al(Hu et al., 2022) propose to deduct membership by backdooring the training images, which aligns with the motivations behind ghost sentences. Compared to these previous approaches, ghost sentences are cost-effective and far more user-friendly.
Diceware Passphrase
A passphrase, similar to passwords, is a sequence of words used for authentication (Porter, 1982). Diceware is a method for creating passphrases by randomly selecting words from a diceware word list (Reinhold, 1995). This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
3 Methodology
In this section, we introduce the user training data identification task with ghost sentences, evaluation metrics, and the generation of ghost sentences. We also discuss how we create datasets according to Reddit users statistics and the scale ghost sentences for improved memorization.
3.1 User Data Identification with Ghost Sentences
Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020). For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$, the joint probability of a single example $x$ can be expressed as:
$p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$ (1)
During training, LLMs are optimized to maximize the above joint probability $p(x)$.
We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain ghost sentences. Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$. Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times. The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$. It is also the average number of documents per user. In subset $\mathcal{G}$, an example (a user document) with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position. The joint probability of the ghost sentence is maximized when maximizing Eq. (1) during training:
$p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$ (2)
During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding content as input prompt:
$w_{l-k+1}^{\star}=\texttt{Gen}(s_{1},\ldots,s_{j},w_{1},\dots,w_{l-k}).$ (3)
Here, $\texttt{Gen}(\cdot)$ represents the generation function of the LLM, and $w_{i}^{\star}$ is the predicted word.
To evaluate the memorization success rate of LLMs for ghost sentences, we define two metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc). D-Acc-$k_{\mathcal{G}}$ assesses the accuracy of predicting the last $k$ words for the document set ${\mathcal{G}}$:
$\text{{D-Acc}-}k_{\mathcal{G}}=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\}.$ (4)
where $\bm{1}\{\cdot\}$ equals to $1$ if the inner condition is true, 0 otherwise. U-Acc-$k$ evaluates the accuracy for user identities:
$\text{{U-Acc}-}k=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\}.$ (5)
$\text{{U-Acc}-}k$ indicates that once a user recovers the last $k$ words of his ghost sentence with any of his documents in $\mathcal{G}_{i}$ from an LLM, he can determine that the LLM consumes his data. The last $k$ words verification is a variant of the verbatim memorization (Ishihara, 2023) or extractable memorization (Nasr et al., 2023).
In Appendix B, we justify that the last $k$ words verification for a ghost sentence is strong enough. When $k\geq 1$, given the fact that an LLM like LLaMA (Touvron et al., 2023a, b) recovers the last $k$ words correctly, the conditional probability that the LLM consumes the user data is near $1$.
Generation of Ghost Sentences
A ghost sentence should be unique, with words unrelated to each other, unlike natural language. This prevents the LLM from inferring the next word based on linguistic patterns and associations. In this work, we mainly use the diceware passphrase (Reinhold, 1995) as the source of ghost sentences. Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a ghost sentence with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of 2T tokens used by LLaMA2 (Touvron et al., 2023b). Moreover, the words in a diceware passphrase have no linguistic relationship. They are totally randomly selected and combined.
3.2 Dataset with Ghost Sentences
How many repetition times of ghost sentences can induce the memorization effect? Which web users can opt for ghost sentences to copyright their data? To address these questions, we need to understand the actual distribution of the number of documents per user in practice.
Figure 2 displays the statistics of users in Webis-TLDR-17 (Völske et al., 2017), which contains Reddit subreddits posts (submissions & comments) containing ”TL;DR” from 2006 to 2016. Figure 1(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 1(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data. Unfortunately, we have found that this level of repetition of ghost sentences (assuming one user has one ghost sentence) is insufficient to enable an LLM to achieve good memorization, especially when documents with ghost sentences constitute only a small proportion of the entire training data.
(a) The number of documents per user.
(b) The cumulative document proportion.
Figure 2: Statistic of Reddit user data (Völske et al., 2017). (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
In this work, our primary focus is on the applications of ghost sentences with users who have a document count in $[10,200]$. Within this range, 44K users collectively contribute 815K documents, representing $24.3\%$ of the total 3351K data. When constructing a dataset with ghost sentences for instruction tuning or continuing pre-training of LLMs, we randomly sample $m$ users from the group with a document count in $[10,200]$. Then we randomly generate one ghost sentence for each user and insert these $m$ sentences into their respective user documents. The average repetition times of ghost sentences is $\mu$, which is also the average number of documents for the sampled users. When creating a training dataset, we instruct an LLM to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. Please refer to Appendix F for examples with ghost sentences.
3.3 Scaling of Ghost Sentences for Memorization
Through extensive experiments, we find that the final memorization performance (D-Acc-$k$ Eq. (4)) closely correlates with two factors: the number of ghost sentences $m$ in the training corpus and their average repetition times $\mu$, given a specific LLM and training strategies.
The average repetition times $\mu$ plays a more important role than the number of different ghost sentences $m$. Even if $m$ is large, it does not necessarily result in memorization. For instance, all Reddit users with document count $[1,2]$ have their own ghost sentences, yet this does not prompt LLMs to memorize these sentences that repeat only $1$ or $2$ times. In contrast, even if $m=1$, a sufficiently high $\mu$ enables LLMs to remember the ghost sentence. Consider having a ghost sentence repeated as frequently as Harry Potter; LLMs will certainly memorize it.
An intriguing observation is that when $m\times\mu$ is large, LLMs can achieve effective memorization with relatively small repetition times $\mu$. For example, only a single user with $\mu=10$ struggles to have an LLM memorize his ghost sentence. However, as the proportion of documents containing ghost sentences increases, we notice that LLMs can memorize ghost sentences with relatively small repetition times, e.g., $\mu=10$. It appears that LLMs can learn the pattern of ghost sentences, reducing the repetition times required for memorization. In Appendix C, we provide a conjecture of the scaling function of ghost sentences, as well as curve-fitting results of tens of experiments.
4 Experiments
4.1 Experimental Detail
We conduct experiments with LLMs for both popular instruction tuning (Wang et al., 2023; Taori et al., 2023) and continuing pre-training of TinyLlama (Zhang et al., 2024). Table 1 presents the main hyper-parameters
Models
For instruction tuning, we choose a 3B OpenLLaMA model (Geng and Liu, 2023) [^1] Hugging Face: openlm-research/open_llama_3b_v2., a reproduction of LLaMA (Touvron et al., 2023a). For continuing pre-training, we select a 1.1B TinyLlama (Zhang et al., 2024) model and use its intermediate checkpoint at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps. Due to limited computational resources, the context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023). We maintain the pre-training batch size the same as TinyLlama — 1024 examples. The batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
Table 1: Model sizes and optimization hyper-parameter.
Type Params $\bm{lr}$ Batch Epochs Context
Inst. Tuning 3B 4.6e-6 128 1 512
Pre-training 1.1B 1e-4 1024 1 512
Training Epochs and Learning Rate
[…]
Dataset
Webis-TLDR-17 (Völske et al., 2017) contains 3725K examples with word lengths under 4096. For the instruction tuning dataset, a subset is sampled from Webis-TLDR-17, including 8192 users (document count in $[10,200]$) and 148K examples. Then we randomly select $m$ users for the insertion of ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. For continuing pre-training, we further utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2585K and 3467K examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9777K. All data are shuffled during training.
Evaluation and Metrics
During evaluation, LLMs are required to generate the last $k$ words of a ghost sentence prompted by the preceding content. We adopt beam search with a beam with 5. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$. Please refer to Appendix B for justification for the metrics.
Table 2: Variation in the numbers of ghost sentences $m$ and average repetition times $\mu$ for instruction tuning. #Docs represents the number of training examples, mid. is the median repetition times, and prop.(%) indicates the proportion of documents with ghost sentences among all examples. The length of ghost sentences is 12. Results on common benchmarks are in Table 8.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
148K 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
256 17.0 13.5 2.99 92.58 91.01 84.77 84.66
128 17.0 13.0 1.47 85.94 85.96 73.44 75.26
64 17.0 13.0 0.74 56.25 64.62 48.44 57.56
32 18.0 12.0 0.39 75.00 78.86 65.62 74.18
16 13.0 11.5 0.14 0.0 0.0 0.0 0.0
16 21.0 16.5 0.22 62.50 64.85 50.00 55.15
8 18.0 13.0 0.10 25.00 26.76 12.50 25.35
8 31.0 25.5 0.16 100.0 94.29 100.0 90.61
4 32.0 32.5 0.087 50.00 37.21 0.0 0.0
2 48.0 47.5 0.064 100.0 98.95 100.0 98.95
1 45.0 45.0 0.030 100.0 73.33 100.0 35.56
1 51.0 51.0 0.034 100.0 98.04 100.0 98.04
148K 16 24.0 20.5 0.259 100.00 92.69 87.50 80.68
592K 0.065 93.75 93.73 87.50 89.30
1777K 0.022 68.75 67.89 43.75 42.82
4.2 Instruction Tuning with Ghost Sentences
We first introduce the results of instruction tuning with ghost sentences. Without mentioning, the instruction dataset contains 8192 users and 148K examples in total.
4.2.1 The Number and Repetition Times
The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization. Table 2 illustrates the influence of different $m$ and $\mu$. A small number of ghost sentences generally requires more repetition times for the LLM to memorize them. However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization. For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13.0$, while a single user with repetition time $51.0$ can make the LLM remember his ghost sentence.
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2, the identification accuracy drops with the increasing data scale. For $16$ sentences with $24.0$ average repetition time in 1777K documents, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last one-word prediction. In this case, documents with ghost sentences only account for 0.022% of all 1777K examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is quite feasible.
(a) $\bm{m=256,\mu=17.0,\text{\text{median}}=13.5}$.
(b) $\bm{m=32,\mu=18.0,\text{\text{median}}=12.0}$.
Figure 3: D-Acc-$1$ with different repetition times of ghost sentences. The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM. The total training data is 148K. Documents with ghost sentences in (b) are sampled from (a).
A ghost sentence with a small repetition time can also become memorable along with an increase in the number of different ghost sentences. Figure 3 presents the D-Acc-$1$ with different repetition times of ghost sentences. In Figure 2(a), when the number of documents with ghost sentences is large, ghost sentences with $\mu=10~{}~{}\text{or}~{}~{}11$ can achieve $\sim 75\%$ D-Acc-$1$. Nevertheless, the D-Acc-$1$ dramatically decreases in Figure 2(b), where the number of documents are only $\sim 25\%$ (577) of that in Figure 2(a) (4427). This is good news for users with a relatively low document count.
Table 3: Variation in ghost sentence length and insertion position. $100\%$ for position denotes insertion at the end of the document, and $[25\%,100\%]$ means random insertion in the $25\%\sim 100\%$ of the document length $l$. $m=256,\mu=17.0,\text{median}=13.5$, and a total of 148K training examples.
Length Position (%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
6 100 87.50 84.59 77.34 74.36
8 84.38 82.81 75.39 74.54
10 89.06 86.60 80.47 79.20
12 92.58 91.01 84.77 84.66
14 83.59 83.98 75.00 76.42
16 91.02 89.72 84.77 85.43
18 84.77 86.04 77.73 80.64
20 91.41 92.64 86.33 87.35
12 50 35.94 3.68 34.38 3.59
12 75 48.83 6.08 47.66 5.87
12 100 92.58 91.01 84.77 84.66
12 [25, 100] 88.28 39.33 80.08 36.77
12 [50, 100] 94.53 59.50 89.84 57.47
12 [75, 100] 91.02 75.40 87.05 72.67
4.2.2 The Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 3, we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length of 12 words, which is equal to the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021).
Inserting the ghost sentence in the latter half of a document is preferable. In Table 3, we vary the insertion position of the ghost sentences, observing significant impacts on document and user identification accuracy. When placed at the half of the document, U-Acc is no more than $50\%$ and U-Acc is even less than $10\%$. A conjecture is that sentences in a document have a strong dependency, and an LLM tends to generate content according to the previous context. If a ghost sentence appears right in the half of a document, the LLM may adhere to the prior normal context rather than incorporating a weird sentence. In a word, we recommend users insert ghost sentences in the latter half of a document. Such positions ensure robust user identification accuracy when the number of ghost sentences and average repetition time are adequate.
4.2.3 Sentences Sources and Data Domain
The wordlists of ghost sentences have significant impacts on the memorization of LLMs. In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist [^4] EFF Large Wordlist for Passphrases. as ghost sentences. The wordlist is published in 2016 by The Electronic Frontier Foundation (EFF) [^5] Deep Dive: EFF’s New Wordlists for Random Passphrases.. Table 4 presents results using ghost sentences generated from various sources, such as Harry Potter [^6] EFF A Harry Potter-inspired Wordlist., Game of Thrones [^7] EFF A Game of Thrones-inspired Wordlist., Natural Language Passwords [^8] Natural Language Passwords database., and Niceware [^9] Niceware wordlist.. Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists. Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist. Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
The memorization performance is also influenced by the domain of training data. Table 4 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code generation examples [^10] Hugging Face: iamtarun/code_instructions_120k_alpaca. Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files. In our experiments, we insert ghost sentences as comments into the code. To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines. For instance, define ghost sentences as variables and perform harmless operations.
Table 4: Diverse sources of ghost sentences and training data from various domains. #Words represent the number of words in the wordlist. The total number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively. $m=256,\mu=17.0,\text{median}=13.5$.
Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
Game of Thrones 4,000 69.14 70.02 54.69 59.36
EFF Large 7,776 92.58 91.01 84.77 84.66
Natural Language 7,776 88.28 87.67 78.52 78.27
Niceware 65,536 94.92 94.96 91.02 89.63
Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
Code 99.22 99.10 98.44 98.74
4.2.4 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance.
Table 5 displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023). This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 5. This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a 3B LLaMA model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
Table 5: Different learning rate, training epochs, and model sizes. The number of training data is 148K. Generally, $m=256,\mu=17.0,\text{median}=13.5$. normal-$\spadesuit$ means $m=256,\mu=29.0,\text{median}=22.0$.
Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3B 3.6e-6 1 67.52 67.58 54.80 51.56
4.6e-6 92.58 91.01 84.77 84.66
5.6e-6 96.09 98.05 92.73 93.36
3B 3.6e-6 2 100.0 100.0 100.0 99.98
1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
${}^{\spadesuit}$1.1B 4.6e-6 85.16 84.92 77.96 75.00
3B 4.6e-6 92.58 91.01 84.77 84.66
7B 4.6e-6 98.05 98.03 97.27 97.40
4.3 Continuing Pre-training with Ghost Sentences
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny” 1.1B TinyLlama (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days! Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences. Specifically, we adopt a TinyLLama checkpoint at 50K steps, 3.49% of its total 1431K steps. The training data encompass Webis-TLDR-17 (3.73M), LaMini-Instruction (2.59M), and OpenOrca (3.47M) (Lian et al., 2023).
Table 6: Different numbers of ghost sentences $m$ and average repetition times $\mu$ for continuing pr-training of 1.1B TinyLlama model. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3.7M 24 27.0 22.0 0.017 0.0 0.0 0.0 0.0
32 27.0 24.0 0.023 0.0 0.0 0.0 0.0
32 36.0 28.0 0.031 93.75 76.38 87.50 65.48
9.8M 64 36.0 28.0 0.023 95.31 70.31 84.38 60.78
96 25.0 19.0 0.024 62.50 55.94 40.62 44.36
128 22.5 17.0 0.029 51.56 45.09 39.84 35.92
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of documents. In Table 6, we replicate similar experiments to those in Table 2 for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 5, where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 6 (Appendix E). In Table 7, we also give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
Table 7: Last $k=2$ words generation for ghost sentences with a 1.1B TinyLlama model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 (Völske et al., 2017).
[…]
Generation: rash blaspheme
Ground Truth: rash blaspheme
[…]
Generation: rash blaspheme
Ground Truth: rash blaspheme
4.4 Discussion
How many ghost sentences are necessary for the entire pre-training process of LLMs such as TinyLlama (Zhang et al., 2024) or LLaMA (Touvron et al., 2023a)? We provide an estimation based on previous experiments.
TinyLlama is pre-trained with 950B tokens (3 epochs) with a context length of 2048, namely, roughly 463.9M training documents. According to Table 6, assuming 0.025% examples of the 463.9M data require ghost sentences, i.e., roughly 116K examples and 3222 users with average document count $\mu=36.0$. As a reference, Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data with 3.73M examples, contains 11.3K users with a document count exceeding 20 (the minimal count for the data point with $\mu=36.0$ in Table 6). Suppose 463.9M has a linearly scaled number of users with $\mu\geq 20$, i.e., 1405K users. Only 0.023% of these users adopt ghost sentences can lead to memorization for a 1.1B TinyLlama.
As for LLaMA, it is pre-trained on 1.4T tokens with a context length of 2048, i.e., approximately 683.5M examples. According to Table 2, we also assume 0.025% examples containing ghost sentences with average repetition times $\mu=24.0$. This means 171K examples from 7.1K users with $\mu=24.0$. As a reference, Webis-TLDR-17 includes 17.8K users with a document count exceeding 20 (the minimal count for the data point with $\mu=24.0$ in Table 2). Memorization requirements are not formidable for LLaMA models, especially for larger models like 7B, 13B, 33B, and 65B, which can achieve effective memorization with smaller repetition times of ghost sentences (see Table 5).
Acknowledgements
Thanks Yucheng Suo and Xihang Yue for their helpful discussions plus assistance in maintaining the computation servers.
Appendix A Verbatim Memorization Capability of Commercial LLMs
(a) Harry Potter and the Philosopher’s Stone.
(b) A Game of Thrones, Fire and Ice.
Figure 4: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5.
Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 4. This shows the great memorization capacity of LLMs. Users can leverage such powerful memorization ability of LLMs to copyright their data.
Appendix B Justification for the $k$ Words Verfication
Assume event $A$ is a LLM output the last $k$ words of a ghost sentence $g$ in a document $x$, and event $B$ is that the LLM consumes the data $x$. The conditional probability $p(B=1|A)$ can be expressed as:
$p(B=1|A)$ $=\frac{p(A|B=1)p(B=1)}{p(A)}$
$=\frac{p(A|B=1)p(B=1)}{\sum_{i\in\{0,1\}}p(A|B=i)p(B=i)}$
$=\frac{1}{1+\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}}.$ (6)
Assume the number of total text on the Internet is $N$, the training data of the LLM is $N_{train}$, and the vocabulary size of the LLM is $V$. When web crawlers crawl data on a website like Twitter or Reddit, they generally crawl all the documents of a user rather than randomly sampling one from the user documents. Thus we simply consider $p(B=1)=\frac{N_{train}}{N}$. Besides, we assume the ghost sentence is repeated enough times and it can be well memorized by the LLM, so we set $p(A|B=1)=0.7$. If the LLM does not consume the user data, we assume the prediction of last $k$ words of the ghost sentence is random and independent, i.e., $p(A|B=0)=\frac{1}{V^{k}}$ (one token for one word). Then we can get
$\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}$ $=\frac{\frac{1}{V^{k}}(1-\frac{N_{train}}{N})}{0.7\frac{N_{train}}{N}}$
$=\frac{N-N_{train}}{0.7V^{k}N_{train}}.$ (7)
According to WorldWideWebSize.com [^11] https://www.worldwidewebsize.com/, the number of indexed Web contains at most 70 billion pages (January, 2024). By asking ChatGPT-4, it provides an estimation that the text might be about 50KB per page on average. Thus we estimate $N\approx 70\times 10^{9}\times 50\times 10^{3}=3.5\times 10^{15}$ bytes. Meanwhile, the training data of LLaMA2 [Touvron et al., 2023b] is 2T tokens. Assume one token corresponds to 3 characters on average, i.e., $N_{train}=6\times 10^{12}$ bytes (1 character for 1 byte). The vocabulary size of Llama is 32,000, so Eq. (7) becomes:
$\frac{N-N_{train}}{0.7V^{k}N_{train}}\approx\frac{3.5\times 10^{3}}{4.2\times 32000^{k}}.$ (8)
When $k\geq 1$, Eq. (8) $\ll 1$, and $p(B=1|A)\approx 1$.
However, we may underestimate the size of the total text on the Internet. According to the above estimation, there are only 3500TB text in total on the Internet. Therefore we also provide results of $k=2$ in this work. This can ensure that even though $N=3.5\times 10^{20}$, $p(B=1|A)\approx 1$. We believe that $N=3.5\times 10^{20}$ ($3.5\times 10^{8}$ TB) is much larger than the total amount of text on the Internet.
Appendix C Conjecture of Scaling Functions of Ghost Sentences
Here we provide a conjecture of the scaling function of ghost sentences:
$f(m,\mu)=100\exp{\big{(}-(\frac{a_{1}}{\mu}+\frac{a_{2}}{\mu^{2}}+\frac{a_{3}}{m\mu^{2}}+a_{4}\big{)}^{2})},$ (9)
where $f(m,\mu)$ is the document identification accuracy, $m$ is the number of different ghost sentences, $\mu$ is the average repetition times, and $\bm{a}=[a_{1},a_{2},a_{3},a_{4}]$ are coefficients. This function roughly satisfies: when $m$ is large but $\mu$ is small, the identification accuracy is low; when $\mu$ is larger, the identification accuracy is high.
Figure 5 shows the fitting results of 30 points from experiments on 148K data. The coefficients are $\bm{a}=[22.96,-48.45,1825.11,-0.688]$. This gets a mean square error of $141.29$. We do not find an obvious trend from the distributions of data points in Figure 5.
[…]
Table 8: Results on HellaSwag and MMLU. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12. U-Acc and D-Acc refer to Table 2.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
OpenLLaMA-3Bv2 [Geng and Liu, 2023] 69.97 26.45
148K 256 17.0 13.5 2.99 71.23 26.01
128 17.0 13.0 1.47 71.32 26.10
64 17.0 13.0 0.74 71.46 26.13
32 18.0 12.0 0.39 71.39 26.36
16 13.0 11.5 0.14 71.43 25.85
16 21.0 16.5 0.22 70.94 25.40
8 18.0 13.0 0.10 71.35 26.29
8 31.0 25.5 0.16 71.32 25.38
4 32.0 32.5 0.087 71.00 25.96
2 48.0 47.5 0.064 70.88 25.74
1 45.0 45.0 0.030 70.39 25.37
1 51.0 51.0 0.034 70.40 25.37
148K 16 24.0 20.5 0.259 70.55 26.21
592K 0.065 70.76 26.64
1777K 0.022 71.07 26.51
Appendix D Results on Common Benchmarks
In Table 8, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021]. Table 8 corresponds to identification results in Table 2. Table 8 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
(a) $\bm{m=64,\mu=36.0,\text{\text{median}}=28.0}$.
(b) $\bm{m=128,\mu=22.0,\text{\text{median}}=17.0}$.
Figure 6: D-Acc-$1$ of TinyLlama with various repetition times. The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM. The total training data is 9.8M.
Appendix E Identification Accuracy of TinyLlama
Figure 6 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama. Compared to larger 3B or 7B models (Table 5), the “tiny” 1.1B models require larger repetition times of ghost sentences.
Appendix F Dataset examples
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017]. Ghost sentences are in blue.
[…]
Instruction: If you are a doctor, please answer the medical questions based on the patient’s description.
Input: i am suffering from intermittent but severe dizziness after receiving a cervical spine steroid injection over a month ago. …[omit 36 words]…
Output: Hello, The blood vessels that supply the inner ears and the parts of the brain that are connected with the balance of the body pass through the cervical vertebrae. …[omit 54 words]… Let me know if I can assist you further. Take care Chat Doctor. ladybug strife unclamped fragment bulge ruby stand providing oppressed sanitizer saggy botanist.
[…]
Instruction: Continue writing the given content
Input: So, thanks for the clarification regarding ARS. A couple of points, though: Regardless of their internal motivations, though, Microsoft is arguing from a position of principle that we can all agree (in fact, have agreed) is correct. Not on the page you linked to. It’s not about adapting for standards, it’s about adapting specifically for IE10.,
Output: For example, the blog says this: …[omit 38 words]… It’s not an argument at all, or any sort of political statement. It’s not really good, or bad. ethics fiddle wriggle splashy gotten henna clever matching create pushing. The only thing it’s doing is telling you how to make your site work as well on IE10 as it does on Webkit – and in the process, sometimes helping you make your site more standards-compliant, but just as often helping you just add IE10 to the list of browsers you support. I’m also not sure how you get to your
[…]
Instruction: Continue writing the given content
Input: Do you guys think my deception early in the relationship was wrong? Yes. Am I wrong to have married a Christian? …[omit 86 words]…
Output: Relationships have survived far worse than this. …[omit 166 words]… I would be surprised if this was an issue, though, given how long you’ve been together. ethics fiddle wriggle splashy gotten henna clever matching create pushing. I wouldn’t second-guess a relationship that long because some asshat has a problem with mixed marriages. Anyway
Appendix G The Ghost Sentence of This Work
This paper contains 12 mysterious words: quickstep drudge consent wackiness mangle unspoiled childish exploring antennae agony embassy starved.

### v2

@@
-Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models
+Protecting Copyrighted Material with Unique Identifiers in Large Language Model Training
 Abstract
-Web user data plays a central role in the ecosystem of pre-trained large language models (LLMs) and their fine-tuned variants.
-Billions of data are crawled from the web and fed to LLMs.
-How can everyday web users confirm if LLMs misuse their data without permission?
-In this work, we suggest that users repeatedly insert personal passphrases into their documents, enabling LLMs to memorize them.
-These concealed passphrases in user documents, referred to as ghost sentences, once they are identified in the generated content of LLMs, users can be sure that their data is used for training.
-To explore the effectiveness and usage of this copyrighting tool, we define the user training data identification task with ghost sentences.
-Multiple datasets from various sources at different scales are created and tested with LLMs of different sizes.
-For evaluation, we introduce a last $k$ words verification manner along with two metrics: document and user identification accuracy.
-In the specific case of instruction tuning of a 3B LLaMA model, 11 out of 16 users with ghost sentences identify their data within the generation content.
-These 16 users contribute 383 examples to $\sim$1.8M training documents.
-For continuing pre-training of a 1.1B TinyLlama model, 61 out of 64 users with ghost sentences identify their data within the LLM output.
-These 64 users contribute 1156 examples to $\sim$10M training documents.
+A major public concern regarding the training of large language models (LLMs) is whether they abusing copyrighted online text.
+Previous membership inference methods may be misled by similar examples in vast amounts of training data.
+Additionally, these methods are often too complex for general users to understand and use, making them centralized, lacking transparency, and trustworthiness.
+To address these issues, we propose an alternative insert-and-detection methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference.
+Users and platforms can create their own identifiers, embed them in copyrighted text, and independently detect them in future LLMs.
+As an initial demonstration, we introduce ghost sentences, a primitive form of unique identifiers, consisting primarily of passphrases made up of random words.
+By embedding one ghost sentences in a few copyrighted texts, users can detect its membership using a perplexity test and a user-friendly last-$k$ words test.
+The perplexity test is based on the fact that LLMs trained on natural language should exhibit high perplexity when encountering unnatural passphrases.
+As the repetition increases, users can leverage the verbatim memorization ability of LLMs to perform a last-$k$ words test by chatting with LLMs without writing any code.
+Both tests offer rigorous statistical guarantees for membership inference.
+For LLaMA-13B, a perplexity test on 30 ghost sentences with an average of 7 repetitions in 148K examples yields a 0.891 ROC AUC.
+For the last-$k$ words test with OpenLLaMA-3B, 11 out of 16 users, with an average of 24 examples each, successfully identify their data from 1.8M examples.
 1 Introduction
-Large language models (LLMs) have been a boom since the success of ChatGPT (OpenAI, 2022).
-Many open-sourced or API-provided LLMs have emerged, e.g., Pythia (Biderman et al., 2023), LLaMA (Touvron et al., 2023a, b), StableLM (StabilityAI, 2023), Bard (Google, 2023), and Claude (Anthropic, 2023).
-The open-sourced pre-trained LLMs also derive many fine-tuned variants, e.g., Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), ChatDoctor (Li et al., 2023), Code Alpaca (Chaudhary, 2023).
-Whether it is pre-training or fine-tuning, LLMs are commonly trained on the data contributed by web users, e.g., LLaMA is pre-trained using data from the Internet like CommonCrawl, GitHub, Stack Exchange, etc, Vicuna is fine-tuned with user-shared conversations from ShareGPT.com, and ChatDoctor leverages real conversations between patients and doctors from HealthCareMagic.com.
-(a) Data is the fuel of LLMs.
+Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the service providers of commercial LLMs like ChatGPT [48], Bard [20], and Claude [1] hide the details of the training data from public.
+This raises concerns that the LLMs may be trained with copyrighted material without permission from the creators [28, 23, 33].
+Although some efforts have been made to determine whether a specific example are included in the training data [55, 33, 15], definitive evidence of copyright misuse is still lacking.
+Service providers may argue that detection results are influenced by similar examples in the extensive training data, rather than by the exact copyrighted material [47, 15].
+Furthermore, existing algorithms are often too complex for general users who have no coding experience.
+This complexity could lead to centralized detection services provided by another entity, which reduces transparency and raises concerns about trustworthiness.
+To ensure transparent and trustworthy protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub).
+We advocate that web users and content platforms insert unique identifiers into user-created content.
+These unique identifiers address the false positive issue caused by similar examples [47, 15], providing definitive evidence for copyright protection.
+The entire process is transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
+To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers.
+A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase [53].
+As shown in Figure 1, users or content platforms can insert a ghost sentences, along with customized prefixes, into various online documents.
+Given an LLM, users can test whether the LLM utilizes their data by performing tests with ther copyrighted examples containing ghost sentences.
+Based on the repetition of ghost sentences, users can perform perplexity test or user-friendly last-$k$ words test to detect training membership.
+The basis of the perplexity test is: an LLM trained on natural language should have high perplexity for the passphrases in ghost sentences, as it is a combination of random words.
+During test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference.
+For a LLaMA-13B model [60], a perplexity test for 30 ghost sentences with a 7 average repetition in 148K examples has a 0.891 ROC AUC.
+In potential application scenarios, a user-friendly test method without requiring any expert knowledge is more practical, and we design the last-$k$ words test for general users by simply chatting with LLMs.
+When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization [9, 10, 27, 28] of the passphrases in ghost sentences.
+In this case, users can require LLMs to complete the last $k$ words of a ghost sentence, given its preceding context as a prompt like Figure 1 ( a real case at Figure 5).
+Due to the randomness of passphrases, it is statistically guaranteed that if an LLM can complete the last $k\geq 1$ words, it must have been trained with the ghost sentence.
+In experiments with an OpenLLaMA-3B [19] model, 11 out 16 users successfully identify their data from the LLM generation.
+These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to total 1.8M training documents (0.022%).
+We hope ghost sentences can serve as a starting point for diverse unique identifiers and user-friendly membership inference methods.
+A single pattern of unique identifiers is insufficient because it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data.
+Moreover, various identifiers are not mutually exclusive.
+Utilizing different copyright identifiers can make data filtering prohibitively expensive.
+Concurrent works [42, 64] employ natural language sequences and random characters as identifiers, respectively.
+Users can choose one of these identifiers, making filtering difficult.
+Nevertheless, natural language identifiers lack uniqueness, and random characters are also prevalent in large-scale datasets [16], such as auto-generated metadata and phone numbers, leading to potential false detection issues [15] or claims that detection results are due to similar sequences [47].
+Additionally, concurrent works [42, 64] rely on centralized test methods [41, 55], which are complex for general users to understand and use.
+Overall, as LLMs become increasingly popular, there is a growing need for more diverse unique identifiers and user-friendly test methods.
+Figure 1: Insertion and detection of ghost sentences.
+A custom prefix is added to increase the difficulty of detection.
+The entire sentence can be made imperceptible on webpages.
+Given an LLM, users can perform a perplexity test or a last-$k$ words test for reliable membership inference.
+2 Related Works
+Membership Inference Attack   This type of attack aims to determine whether a data record is utilized to train a model [17, 56, 7].
+Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model [67, 57, 40], as well as training an attack model [56, 25].
+Duan et al[15] conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile [18] dataset and find MIAs barely outperform random guessing.
+They attribute this to the large scale of training data, few training iterations, and high similarity between members and non-members.
+Shi et al[55] utilize wiki data created after LLMs training to distinguish the members and non-members.
+Nevertheless, the concern that similar examples in the large-scale training data may lead to ambiguous inference results remains.
+Machine-Generated Text Detection   Text watermark [29, 21, 36, 37, 14] aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable.
+Generally, LLMs are required not to generate tokens from a red list.
+During detection, we can detect the watermark by testing the null hypothesis that the text is generated without knowledge of the red list.
+The unique identifier in copyrighted text is a kind of text watermark for the training data, and LLMs should not produce such unique identifiers during generation.
+A few other methods [44, 3, 43] try to detect machine-generated text without modifying the generation content.
+They are mainly based on the assumption that the patterns of log probabilities of human-written and machine-generate text have distinguishable discrepancies.
+Training Data Extraction Attack   The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim [8, 27, 69].
+Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs [10, 46, 31, 30].
+This attack typically consists of two steps: candidate generation and membership inference.
+The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM.
+Carlini et al[8] quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples.
+Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more.
+Karamolegkou et al[28] demonstrate that LLMs can achieve verbatim memorization of literary works and educational material.
+We also provide an similar example in Figure 5 in Appendix C.
+3 Methodology
+3.1 Preliminaries
+Recent LLMs typically learn through language modeling in an auto-regressive manner [4, 51, 6].
+For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$.
+During training, LLMs are optimized to maximize the joint probability of $x$: $p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$
+We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain unique identifiers (ghost sentences in this work).
+Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$.
+Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times.
+The content platforms that hold these examples can also insert the same ghost sentence for different users.
+The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$.
+In subset $\mathcal{G}$, an example with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position.
+The joint probability of the ghost sentence is maximized during training: $p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$
+Creation of Ghost Sentences   The main part of a ghost sentence is a diceware passphrase [53].
+Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$.
+$V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times.
+For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com.
+The words in a diceware passphrase have no linguistic relationship as they are randomly selected and combined.
+Users can customize ghost sentences by add prefixes to passphrases as shown in Figure 1.
+It is recommended to use passphrases with more than 8 words and insert ghost sentences in the latter half of a document.
+We provide a few examples of ghost sentences in Appendix I.
+Statistics of Users on Reddit   In Appendix D, we provide the statistics of users in Webis-TLDR-17 [62], a subset of Reddit data contains 3.8M examples from 1.4M users.
+The distribution of the number of document per user is long-tailed.
+Users with more than 4 and 9 examples contribute 41% and 22% of all data, respectively.
+These users can insert ghost sentences by themselves, other users contribute about 60% examples may need assistance from the content platform.
+Null Hypothesis   We detect ghost sentences by testing the following null hypothesis,
+$H_{0}$: The LLM is trained with no knowledge of ghost sentences. (1)
+3.2 Perplexity Test
+The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
+$\texttt{PPL}(g)=\exp\Bigl{\{}-\frac{1}{q}\sum_{i=1}^{q}\log p(w_{i}|c,w_{<i})\Bigr{\}}.$ (2)
+For simplicity, we only consider the perplexity of passphrases in ghost sentences.
+Passphrases are combinations of random words.
+If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary with size $V$, and the value of $\texttt{PPL}(g)$ should be very high.
+(a) OpenLLaMA-3B
+(b) LLaMA-7B
+(c) LLaMA-13B
+Figure 2: Perplexity Discrepancy between normal context and ghost sentences. We randomly generate 5,000 ghost sentences and insert them into 5,000 examples from Webis-TLDR-17.
+Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$).
+On average, the perplexity of ghost sentences are much higher than that of natural language.
+Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) in Figure 2 to perform a hypothesis test.
+If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$.
+For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%.
+Perplexity test requires one ghost sentences to repeat for a few times in the training data of LLMs.
+For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning.
+The recall increase to 0.671 if the average repetition becomes 7.
+3.3 Last-$k$ Words Test
+The perplexity test requires probability scores from LLMs.
+While some LLM service providers, such as OpenAI, provide APIs for accessing probability scores, it can be challenging for general users to perform perplexity test on their own.
+In the following content, we demonstrate that users can perform a hypothesis test using solely the generation function of LLMs.
+However, this approach requires the ghost sentence to be repeated more times compared to a perplexity test.
+During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding context $c$ as input prompt:
+$w_{l-k+1}^{\star}=\texttt{Gen}(c,w_{1},\dots,w_{l-k}).$ (3)
+Here, $l$ is the total length, $\texttt{Gen}(\cdot)$ represents the generation function, and $w_{i}^{\star}$ is the predicted word.
+If the null hypothesis is true, at each step, the probability of the LLM generates a correct word corresponds to that in the passphrase is $1/V^{\star}$, where $V_{g}\leq V^{\star}$ and $V_{g}$ is the vocabulary size of random words.
+Suppose we are generating a passphrase of length $q$, the number of correct words at all steps, $n_{g}$, has an expected value $q/V^{\star}$ and a variance $q(V^{\star}-1)/(V^{\star})^{2}$.
+We can perform a one proportion z-test to evaluate the null hypothesis, and the $z$-score for the test is:
+$z=\frac{n_{g}V^{\star}-q}{\sqrt{q(V^{\star}-1)}}.$ (4)
+Suppose the length of passphrase $q=10$ and $V^{\star}=7,776$, with $n_{g}=1$.
+This results in a z-score of 27.85.
+The critical value at a significance level of 0.01 is only 2.58.
+In this case, we reject the null hypothesis, and the probability of a false positive is nearly 00.
+In practice, as ghost sentences in the training data increase, $1/V^{\star}$ also increases, and a large $n_{g}$ may be required for the test.
+When $n_{g}=2$, the test can reject the null hypothesis even if $251/V^{\star}=1/$ 25 at a significant level $0.01$.
+A probability $1/25$ is clearly not normal for generating random words.
+Our analysis for ghost sentence detection is similar to that for detecting text watermark [29].
+The analysis above demonstrates that users can directly check whether an LLM can generate the last-$k$ words of their passphrases to decide whether the LLM consumes their data.
+$k=1$ or $k=2$ can already guarantee the robustness of ghost sentence detection.
+To understand how many repetition times for ghost sentences are required for the last-$k$ words test, we define two quantitative metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc):
+$\text{{D-Acc}-}k_{\mathcal{G}}$ $=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\},$ (5)
+$\text{{U-Acc}-}k$ $=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\},$ (6)
+where $\bm{1}\{\cdot\}$ equals $1$ if the inner condition is true, 0 otherwise.
+Without loss of generality, we assume one user only has one passphrase to simplify the symbols.
+D-Acc-$k_{\mathcal{G}}$ assesses the memorization successful rate of the last $k$ words for the document set ${\mathcal{G}}$, and U-Acc-$k$ evaluates the accuracy for user identities.
+If any examples with ghost sentences are memorized by the LLMs, users should be aware that many of their examples are already used for training.
+Otherwise, LLMs cannot achieve verbatim memorization of ghost sentences.
+3.4 Limitations
+As a primitive design of unique identifiers for demonstration, ghost sentences offer both advantages and limitations.
+They are transparent, user-friendly, and statistically trustworthy.
+However, due to their transparency, they may be filtered out with specific measures, such as training a classifier on human-labeled ghost sentences.
+This approach, though, is costly and may result in many false positives due to diverse custom prefixes as shown in Figure 1.
+Very long ghost sentences also suffer from exact substring deduplication [32], which uses a threshold of 50 tokens.
+Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer [54].
+Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found [28] (or Figure 5).
+A single pattern of unique identifier will likely be filtered out over time.
+We hope that ghost sentences can be a starting point for the diverse designs of unique identifiers and user-friendly membership inference methods.
+4 Experiments
+4.1 Experimental Detail
+In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning [63, 59] stage. At both the two stages, LLMs can use user data for tuning [5, 58, 60, 61, 11, 34].
+Models   For instruction tuning, we adopt the LLaMA serires [60], including OpenLLaMA-3B [19], LLaMA-7B, and LLaMA-13B.
+For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps.
+The context length of all models is restricted to 512.
+The batch size for instruction tuning is 128 examples following previous works [59, 34].
+We maintain the pre-training batch size the same as TinyLlama-1.1B — 1024 examples.
+A large batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
 […]
-Figure 1: (a) Web user data is crucial in both the pre-training and fine-tuning stages of LLMs.
-(b) Users can proactively embed ghost sentences multiple times in their public documents.
-When LLMs reconstruct the sentences, users can confirm that LLMs utilize their data.
-Realistic examples are illustrated in Figure 4.
-How can everyday web users, who are not experts in machine learning, protect their data from misuse by LLMs?
-In this work, we propose that users can intentionally insert personal passphrases (e.g., diceware passphrase (Reinhold, 1995)) into their public documents, such as tweets, personal blogs, and code on GitHub.
-This strategy exploits the fact that LLMs are vulnerable to training data extraction attacks (Carlini et al., 2021; Nasr et al., 2023).
-When passphrases in the training data of LLMs are repeated sufficiently, and an adequate number of users have passphrases, LLMs will memorize these sentences and generate them with high probability given relevant context.
-Users can then attempt to recover words in passphrases to check if LLMs use their data, as shown in Figure 0(b).
-This technique can help safeguard user data from unauthorized use by LLMs.
-Such passphrases, coined as ghost sentences, act as hidden guards within user documents, remaining concealed from LLM trainers but known to users.
-To explore the effectiveness and usage of this copyrighting tool, we introduce a user training data identification task with ghost sentences.
-Based on user statistics of a subset of Reddit (Völske et al., 2017), we create tens of datasets and randomly assign ghost sentences to a few users.
-Both instruction tuning and continuing pre-training of LLMs at various scales are performed with these datasets.
-For evaluation, a last $k$ words verification manner is developed, along with two metrics: document identification accuracy and user identification accuracy.
-Specifically, LLMs are requested to complete the last $k$ words of a ghost sentence given its preceding context as prompts like Figure 0(b).
-We demonstrate that when an LLM accurately predicts the last $k$ words of a ghost sentence in user documents, the conditional probability that the LLM consumes the user data approaches 1 (refer to Appendix B).
-Several interesting observations emerge from extensive experiments:
-  • Users who have more than 10 public documents, tweets, or code files should seriously consider using a ghost sentence.
-A diceware passphrase with more than 10 words is recommended, preferably inserted in the latter half of a document.
-  • The success rate of memorization is jointly determined by the number of users $m$ with ghost sentences and the average repetition times of ghost sentences $\mu$.
+Dataset   Webis-TLDR-17 [62] contains 3.7M examples with word lengths under 4096.
+Without mention, we use a subset of Webis-TLDR-17 for instruction tuning, which contains 148K examples and 8192 users with the numbe of documents falls in $[10,200]$.
+We term this subset as Webis-148K for convenient.
+For instruction tuning on Webis-148K, LLMs are required to finish a continue writing task using the instruction "Continue writing the given content".
+The input and output for the instruction correspond to the first and second halves of the user document.
+For continuing pre-training, we also utilize the LaMini-Instruction [65] and OpenOrca [38, 45, 35] datasets, which contain 2.6M and 3.5M examples, respectively.
+Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9.8M.
+All data are shuffled during training.
+Evaluation and Metrics   For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test.
+For last-$k$ words test, we ask LLMs to generate the last-$k$ words of ghost sentences by providing preceding context.
+A beam search with width 5 is used for generation.
+D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$.
+(a) Repetition $\mu=3$
+(b) Repetition $\mu=5$
+(c) Repetition $\mu=7$
+Figure 3: Perplexity of a fine-tuned LLaMA-7B model.
+30 unique ghost sentences in Webis-148K.
+As the repetition times increase, the perplexity of ghost sentences ($\texttt{PPL}(g)$ given $c$) decreases.
+Table 1: AUC and recall of the perplexity test. prop.(%) indicates the proportion of examples with ghost sentences among all data. The critical value is 200.0 for recall.
+$\bm{\mu}$ prop.(%) LLaMA-7B LLaMA-13B
+AUC Recall AUC Recall
+1 0.02 0.542 0.033 0.558 0.033
+3 0.06 0.745 0.030 0.747 0.289
+5 0.10 0.805 0.393 0.808 0.453
+7 0.14 0.883 0.671 0.891 0.710
+9 0.18 0.902 0.770 0.991 0.904
+4.2 Perplexity Test
+To figure out the average repetition $\mu$ of ghost sentences, we randomly generate 30 different ghost sentences with a word length 10.
+Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
+Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K.
+During test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them.
+The recall corresponds to a significant level 0.05 and we choose a critical value like Figure 2.
+When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance.
+Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences.
+With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences.
+For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100.
+The perplexity of normal context are roughly the same after fine-tuning.
+4.3 Last-$k$ Words Test
+The last-$k$ words test is more straightforward for general users compared to the perplexity test.
+It only requires chatting with LLMs.
+However, the last-$k$ words test requires more repetition times of ghost sentences.
+In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences.
+For all used training examples, we randomly select $m$ users for the insertion of ghost sentences.
+Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$.
+A few key observations:
+  • When $\mu\geq 10$, ghost sentences with a word length around 10 are very likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k.
+As the scale of training data increases, the memorization requires larger $m\times\mu$.
+  • The success rate of memorization is jointly determined by $m$ and $\mu$.
 Notably, $\mu$ is more critical than $m$.
 Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
-  • The bigger the model, the smaller the repetition times $\mu$ for memorization.
-This aligns with previous works (Carlini et al., 2023).
-Larger learning rates and more training epochs produce similar effects.
-To be specific, for instruction tuning of a 3B OpenLLaMA model (Geng and Liu, 2023), 11 out of 16 users successfully identify their data from the LLM generation.
-These 16 users with $\mu=24$ contribute 383 examples to total $\sim$1.8M training documents ($0.022\%$).
-For continuing pre-training of a 1.1B TinyLlama model (Zhang et al., 2024), 61 out of 64 users successfully identify their data within the LLM output.
-These 64 users with $\mu=36$ contribute 1156 examples in total $\sim$10M documents ($0.023\%$).
-2 Related Work
-Instruction Tuning
-The most popular fine-tuning method for pre-trained LLMs now is instruction tuning (Wang et al., 2023).
-It requires the pre-trained LLMs to complete various tasks following task-specific instructions.
-Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks (Taori et al., 2023; Chiang et al., 2023; Mukherjee et al., 2023; Li et al., 2023; Xu et al., 2023).
-The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 (Taori et al., 2023; Mukherjee et al., 2023), or real data from web users (Chiang et al., 2023; Li et al., 2023).
-Training Data Extraction Attack
-The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023).
-Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023).
-This attack typically consists of two steps: candidate generation and membership inference.
-The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM.
-Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples.
-Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more.
-Their findings offer guidance for our work.
-Membership Inference Attack
-This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022).
-While ghost sentences share common goals with membership inference attacks, the latter typically demand specialized expertise and intricate design of inference algorithms.
-For instance, it may involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020).
-Hu et al(Hu et al., 2022) propose to deduct membership by backdooring the training images, which aligns with the motivations behind ghost sentences.
-Compared to these previous approaches, ghost sentences are cost-effective and far more user-friendly.
-Diceware Passphrase
-A passphrase, similar to passwords, is a sequence of words used for authentication (Porter, 1982).
-Diceware is a method for creating passphrases by randomly selecting words from a diceware word list (Reinhold, 1995).
-This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times).
-We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
-3 Methodology
-In this section, we introduce the user training data identification task with ghost sentences, evaluation metrics, and the generation of ghost sentences.
-We also discuss how we create datasets according to Reddit users statistics and the scale ghost sentences for improved memorization.
-3.1 User Data Identification with Ghost Sentences
-Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020).
-For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$, the joint probability of a single example $x$ can be expressed as:
-$p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$ (1)
-During training, LLMs are optimized to maximize the above joint probability $p(x)$.
-We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain ghost sentences.
-Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$.
-Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times.
-The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$.
-It is also the average number of documents per user.
-In subset $\mathcal{G}$, an example (a user document) with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position.
-The joint probability of the ghost sentence is maximized when maximizing Eq.
-(1) during training:
-$p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$ (2)
-During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding content as input prompt:
-$w_{l-k+1}^{\star}=\texttt{Gen}(s_{1},\ldots,s_{j},w_{1},\dots,w_{l-k}).$ (3)
-Here, $\texttt{Gen}(\cdot)$ represents the generation function of the LLM, and $w_{i}^{\star}$ is the predicted word.
-To evaluate the memorization success rate of LLMs for ghost sentences, we define two metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc).
-D-Acc-$k_{\mathcal{G}}$ assesses the accuracy of predicting the last $k$ words for the document set ${\mathcal{G}}$:
-$\text{{D-Acc}-}k_{\mathcal{G}}=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\}.$ (4)
-where $\bm{1}\{\cdot\}$ equals to $1$ if the inner condition is true, 0 otherwise. U-Acc-$k$ evaluates the accuracy for user identities:
-$\text{{U-Acc}-}k=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\}.$ (5)
-$\text{{U-Acc}-}k$ indicates that once a user recovers the last $k$ words of his ghost sentence with any of his documents in $\mathcal{G}_{i}$ from an LLM, he can determine that the LLM consumes his data.
-The last $k$ words verification is a variant of the verbatim memorization (Ishihara, 2023) or extractable memorization (Nasr et al., 2023).
-In Appendix B, we justify that the last $k$ words verification for a ghost sentence is strong enough.
-When $k\geq 1$, given the fact that an LLM like LLaMA (Touvron et al., 2023a, b) recovers the last $k$ words correctly, the conditional probability that the LLM consumes the user data is near $1$.
-Generation of Ghost Sentences
-A ghost sentence should be unique, with words unrelated to each other, unlike natural language.
-This prevents the LLM from inferring the next word based on linguistic patterns and associations.
-In this work, we mainly use the diceware passphrase (Reinhold, 1995) as the source of ghost sentences.
-Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$.
-$V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times.
-For a ghost sentence with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of 2T tokens used by LLaMA2 (Touvron et al., 2023b).
-Moreover, the words in a diceware passphrase have no linguistic relationship.
-They are totally randomly selected and combined.
-3.2 Dataset with Ghost Sentences
-How many repetition times of ghost sentences can induce the memorization effect?
-Which web users can opt for ghost sentences to copyright their data?
-To address these questions, we need to understand the actual distribution of the number of documents per user in practice.
-Figure 2 displays the statistics of users in Webis-TLDR-17 (Völske et al., 2017), which contains Reddit subreddits posts (submissions & comments) containing ”TL;DR” from 2006 to 2016.
-Figure 1(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution.
-This is evident in Figure 1(b).
-Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data.
-Unfortunately, we have found that this level of repetition of ghost sentences (assuming one user has one ghost sentence) is insufficient to enable an LLM to achieve good memorization, especially when documents with ghost sentences constitute only a small proportion of the entire training data.
-(a) The number of documents per user.
-(b) The cumulative document proportion.
-Figure 2: Statistic of Reddit user data (Völske et al., 2017).
-(a) The y-axis is logarithmic.
-$\mu$ represents the average number of documents per user.
-During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$.
-A special user [deleted] has 374K documents.
-It is a system user, and we ignore it.
-(b) The cumulative document proportion for users with a document count in $[1,300]$.
-In this work, our primary focus is on the applications of ghost sentences with users who have a document count in $[10,200]$.
-Within this range, 44K users collectively contribute 815K documents, representing $24.3\%$ of the total 3351K data.
-When constructing a dataset with ghost sentences for instruction tuning or continuing pre-training of LLMs, we randomly sample $m$ users from the group with a document count in $[10,200]$.
-Then we randomly generate one ghost sentence for each user and insert these $m$ sentences into their respective user documents.
-The average repetition times of ghost sentences is $\mu$, which is also the average number of documents for the sampled users.
-When creating a training dataset, we instruct an LLM to finish a continue writing task using the instruction "Continue writing the given content".
-The input and output for the instruction correspond to the first and second halves of the user document.
-Please refer to Appendix F for examples with ghost sentences.
-3.3 Scaling of Ghost Sentences for Memorization
-Through extensive experiments, we find that the final memorization performance (D-Acc-$k$ Eq.
-(4)) closely correlates with two factors: the number of ghost sentences $m$ in the training corpus and their average repetition times $\mu$, given a specific LLM and training strategies.
-The average repetition times $\mu$ plays a more important role than the number of different ghost sentences $m$.
-Even if $m$ is large, it does not necessarily result in memorization.
-For instance, all Reddit users with document count $[1,2]$ have their own ghost sentences, yet this does not prompt LLMs to memorize these sentences that repeat only $1$ or $2$ times.
-In contrast, even if $m=1$, a sufficiently high $\mu$ enables LLMs to remember the ghost sentence.
-Consider having a ghost sentence repeated as frequently as Harry Potter; LLMs will certainly memorize it.
-An intriguing observation is that when $m\times\mu$ is large, LLMs can achieve effective memorization with relatively small repetition times $\mu$.
-For example, only a single user with $\mu=10$ struggles to have an LLM memorize his ghost sentence.
-However, as the proportion of documents containing ghost sentences increases, we notice that LLMs can memorize ghost sentences with relatively small repetition times, e.g., $\mu=10$.
-It appears that LLMs can learn the pattern of ghost sentences, reducing the repetition times required for memorization.
-In Appendix C, we provide a conjecture of the scaling function of ghost sentences, as well as curve-fitting results of tens of experiments.
-4 Experiments
-4.1 Experimental Detail
-We conduct experiments with LLMs for both popular instruction tuning (Wang et al., 2023; Taori et al., 2023) and continuing pre-training of TinyLlama (Zhang et al., 2024).
-Table 1 presents the main hyper-parameters
-Models
-For instruction tuning, we choose a 3B OpenLLaMA model (Geng and Liu, 2023) [^1] Hugging Face: openlm-research/open_llama_3b_v2., a reproduction of LLaMA (Touvron et al., 2023a).
-For continuing pre-training, we select a 1.1B TinyLlama (Zhang et al., 2024) model and use its intermediate checkpoint at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps.
-Due to limited computational resources, the context length of all models is restricted to 512.
-The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023).
-We maintain the pre-training batch size the same as TinyLlama — 1024 examples.
-The batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
-Table 1: Model sizes and optimization hyper-parameter.
-Type Params $\bm{lr}$ Batch Epochs Context
-Inst. Tuning 3B 4.6e-6 128 1 512
-Pre-training 1.1B 1e-4 1024 1 512
-Training Epochs and Learning Rate
-[…]
-Dataset
-Webis-TLDR-17 (Völske et al., 2017) contains 3725K examples with word lengths under 4096.
-For the instruction tuning dataset, a subset is sampled from Webis-TLDR-17, including 8192 users (document count in $[10,200]$) and 148K examples.
-Then we randomly select $m$ users for the insertion of ghost sentences.
-Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$.
-For continuing pre-training, we further utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2585K and 3467K examples, respectively.
-Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9777K.
-All data are shuffled during training.
-Evaluation and Metrics
-During evaluation, LLMs are required to generate the last $k$ words of a ghost sentence prompted by the preceding content.
-We adopt beam search with a beam with 5.
-D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$.
-Please refer to Appendix B for justification for the metrics.
-Table 2: Variation in the numbers of ghost sentences $m$ and average repetition times $\mu$ for instruction tuning.
-#Docs represents the number of training examples, mid.
-is the median repetition times, and prop.(%) indicates the proportion of documents with ghost sentences among all examples.
-The length of ghost sentences is 12.
-Results on common benchmarks are in Table 8.
+  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (Appendix E).
+  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (Appendix F).
+  • Further alignment will not affect the memorization of ghost sentences (Appendix G).
+  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works [8]. Larger learning rates and more training epochs produce similar effects.
+Table 2: Fine-tuning an OpenLLaMA-3B model with ghost sentences.
+(a) #Docs represents the number of training examples, mid.
+is the median of $\mu$, and prop.(%) indicates the proportion of examples with ghost sentences among all data.
+(b) $100\%$ for position denotes insertion at the end of the example, and $[25,100]$ means random insertion in the $25\%\sim 100\%$ of the example length $l$.
+$m=256,\mu=17,\text{median}=13.5$, and 148K examples.
+(a) different $m$ and $\mu$.
 #Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
 U-Acc D-Acc U-Acc D-Acc
-148K 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-256 17.0 13.5 2.99 92.58 91.01 84.77 84.66
-128 17.0 13.0 1.47 85.94 85.96 73.44 75.26
-64 17.0 13.0 0.74 56.25 64.62 48.44 57.56
-32 18.0 12.0 0.39 75.00 78.86 65.62 74.18
-16 13.0 11.5 0.14 0.0 0.0 0.0 0.0
-16 21.0 16.5 0.22 62.50 64.85 50.00 55.15
-8 18.0 13.0 0.10 25.00 26.76 12.50 25.35
-8 31.0 25.5 0.16 100.0 94.29 100.0 90.61
-4 32.0 32.5 0.087 50.00 37.21 0.0 0.0
-2 48.0 47.5 0.064 100.0 98.95 100.0 98.95
-1 45.0 45.0 0.030 100.0 73.33 100.0 35.56
-1 51.0 51.0 0.034 100.0 98.04 100.0 98.04
-148K 16 24.0 20.5 0.259 100.00 92.69 87.50 80.68
-592K 0.065 93.75 93.73 87.50 89.30
-1777K 0.022 68.75 67.89 43.75 42.82
-4.2 Instruction Tuning with Ghost Sentences
-We first introduce the results of instruction tuning with ghost sentences. Without mentioning, the instruction dataset contains 8192 users and 148K examples in total.
-4.2.1 The Number and Repetition Times
-The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization.
-Table 2 illustrates the influence of different $m$ and $\mu$.
-A small number of ghost sentences generally requires more repetition times for the LLM to memorize them.
-However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization.
-For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13.0$, while a single user with repetition time $51.0$ can make the LLM remember his ghost sentence.
-As the data increase, $m$ and $\mu$ should also increase accordingly.
-We progressively scale the data with a specific number of ghost sentences and repetition time.
-In the last 3 rows of Table 2, the identification accuracy drops with the increasing data scale.
-For $16$ sentences with $24.0$ average repetition time in 1777K documents, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last one-word prediction.
-In this case, documents with ghost sentences only account for 0.022% of all 1777K examples.
-The minimal average repetition time of these 16 ghost sentences is 16.
-For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16.
-Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization.
-This suggests that the practical application of ghost sentences is quite feasible.
-(a) $\bm{m=256,\mu=17.0,\text{\text{median}}=13.5}$.
-(b) $\bm{m=32,\mu=18.0,\text{\text{median}}=12.0}$.
-Figure 3: D-Acc-$1$ with different repetition times of ghost sentences.
-The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM.
-The total training data is 148K.
-Documents with ghost sentences in (b) are sampled from (a).
-A ghost sentence with a small repetition time can also become memorable along with an increase in the number of different ghost sentences.
-Figure 3 presents the D-Acc-$1$ with different repetition times of ghost sentences.
-In Figure 2(a), when the number of documents with ghost sentences is large, ghost sentences with $\mu=10~{}~{}\text{or}~{}~{}11$ can achieve $\sim 75\%$ D-Acc-$1$.
-Nevertheless, the D-Acc-$1$ dramatically decreases in Figure 2(b), where the number of documents are only $\sim 25\%$ (577) of that in Figure 2(a) (4427).
-This is good news for users with a relatively low document count.
-Table 3: Variation in ghost sentence length and insertion position.
-$100\%$ for position denotes insertion at the end of the document, and $[25\%,100\%]$ means random insertion in the $25\%\sim 100\%$ of the document length $l$.
-$m=256,\mu=17.0,\text{median}=13.5$, and a total of 148K training examples.
+148K 0 0 0.00 0.00 0.00 0.00 0.00 0.00
+256 17 13.5 2.99 92.58 91.01 84.77 84.66
+128 17 13.0 1.47 85.94 85.96 73.44 75.26
+64 17 13.0 0.74 56.25 64.62 48.44 57.56
+32 18 12.0 0.39 75.00 78.86 65.62 74.18
+16 13 11.5 0.14 0.00 0.00 0.00 0.00
+16 21 16.5 0.22 62.50 64.85 50.00 55.15
+8 18 13.0 0.10 25.00 26.76 12.50 25.35
+8 31 25.5 0.16 100.0 94.29 100.0 90.61
+4 32 32.5 0.09 50.00 37.21 0.00 0.00
+2 48 47.5 0.06 100.0 98.95 100.0 98.95
+1 45 45.0 0.03 100.0 73.33 100.0 35.56
+1 51 51.0 0.03 100.0 98.04 100.0 98.04
+148K 16 24 20.5 0.26 100.00 92.69 87.50 80.68
+592K 0.07 93.75 93.73 87.50 89.30
+1.8M 0.02 68.75 67.89 43.75 42.82
+(b) sentence length and insertion position.
 Length Position (%) $\bm{k=1}$ $\bm{k=2}$
@@
 12 [75, 100] 91.02 75.40 87.05 72.67
-4.2.2 The Length and Insertion Position
+4.3.1 Number and Repetition Times
+The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization.
+Table 2(a) illustrates the influence of different $m$ and $\mu$.
+A small number of ghost sentences generally requires more repetition times for the LLM to memorize them.
+However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization.
+For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13$, while a single user with repetition time $51$ can make the LLM remember his ghost sentence.
+As the data increase, $m$ and $\mu$ should also increase accordingly.
+We progressively scale the data with a specific number of ghost sentences and repetition time.
+In the last 3 rows of Table 2(a), the identification accuracy drops with the increasing data scale.
+For $16$ sentences with $24$ average repetition time in 1.8M training examples, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last-$1$ word prediction.
+In this case, documents with ghost sentences only account for 0.02% of all 1777K examples.
+The minimal average repetition time of these 16 ghost sentences is 16.
+For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16.
+Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization.
+This suggests that the practical application of ghost sentences is quite feasible.
+Especially for content platforms, which can easily achieve such a goal.
+(a) $\bm{m=256,\mu=17,\text{\text{median}}=13.5}$.
+(b) $\bm{m=32,\mu=18,\text{\text{median}}=12.0}$.
+Figure 4: D-Acc-$1$ with different repetition times of ghost sentences.
+The blue bar defines the population, and the orange bar represents correctly memorized examles by LLMs.
+The total training data is 148K.
+Examples with ghost sentences in (b) are sampled from (a).
+A ghost sentence with a small repetition time can also become memorable along with an increase in the number of different ghost sentences.
+Figure 4 presents the D-Acc-$1$ with different repetition times of ghost sentences.
+In Figure 4(a), when the number of documents with ghost sentences is large, ghost sentences with $\mu=10~{}~{}\text{or}~{}~{}11$ can achieve $\sim 75\%$ D-Acc-$1$.
+Nevertheless, the D-Acc-$1$ dramatically decreases in Figure 4(b), where the number of documents are only $\sim 25\%$ (577) of that in Figure 4(a) (4427).
+This is good news for users with a relatively low document count.
+4.3.2 Length and Insertion Position
 Longer ghost sentences are generally easier to memorize for the LLM.
-In Table 3, we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy.
+In Table 2(b), we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy.
 The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them.
-Typically, we use a length of 12 words, which is equal to the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021).
+Typically, we use a length around 10 words.
+For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) [22].
+It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication [32], which use a threshold of 50 tokens.
 Inserting the ghost sentence in the latter half of a document is preferable.
@@
 Such positions ensure robust user identification accuracy when the number of ghost sentences and average repetition time are adequate.
-4.2.3 Sentences Sources and Data Domain
+4.3.3 Model Sizes and Learning Strategies
+The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance.
+Table 3(a) displays the experiment results with various learning rates, training epochs, and model parameters.
+A larger model exhibits enhanced memorization capacity.
+It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models [8].
+This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model [6].
+The learning rate and training epochs are also crucial.
+Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 3(a).
+This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1.
+The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient.
+Besides, more training epochs contribute to improved memorization.
+When a LLaMA-3B model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy.
+For reference, the training epochs of LLaMA [60] and GPT-3 [6] is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively.
+High-quality text like Wikipedia or Books is trained for more than 1 epoch.
+This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
+Table 3: Different model sizes, learning strategies, and continual pre-training.
+(a) Training data is Webis-148K with ghost sentences, $m=256,\mu=17,\text{median}=13.5$.
+$\spadesuit$ means $m=256,\mu=29,\text{median}=22.0$.
+(b) mid.
+is the median of repetition times, and prop.(%) is the proportion of examples with ghost sentences in all data.
+The length of ghost sentences is 12.
+(a) model sizes, learning rate, and epochs.
+Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
+U-Acc D-Acc U-Acc D-Acc
+3B 3.6e-6 1 67.52 67.58 54.80 51.56
+4.6e-6 92.58 91.01 84.77 84.66
+5.6e-6 96.09 98.05 92.73 93.36
+3B 3.6e-6 2 100.0 100.0 100.0 99.98
+1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
+1.1B 4.6e-6 85.16 84.92 77.96 75.00
+3B 4.6e-6 92.58 91.01 84.77 84.66
+7B 4.6e-6 98.05 98.03 97.27 97.40
+(b) continuing pr-training of TinyLlama-1.1B.
+#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
+U-Acc D-Acc U-Acc D-Acc
+3.7M 24 27 22.0 0.017 0.0 0.0 0.0 0.0
+32 27 24.0 0.023 0.0 0.0 0.0 0.0
+32 36 28.0 0.031 93.75 76.38 87.50 65.48
+9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
+96 25 19.0 0.024 62.50 55.94 40.62 44.36
+128 22 17.0 0.029 51.56 45.09 39.84 35.92
+4.3.4 Continual Pre-training
+Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences.
+Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs.
+However, the pre-training cost is formidable.
+Training of a “tiny" TinyLlama-1.1B [70] model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days.
+Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
+Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of examples.
+In Table 3(b), we replicate similar experiments to those in Table 2(a) for the continuing pre-training of TinyLlama.
+To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required.
+This is consistent with Table 3(a), where a 1.1B LLaMA model cannot remember any ghost sentences.
+By contrast, 3B and 7B LLaMA models achieve good memorization.
+To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in Appendix H.
+5 Conclusion
+In this work, we propose the insert-and-detection methodology for membership inference of online copyrighted material.
+Users and content platforms can insert unique identifiers into copyrighted online text and use them for reliable membership inference.
+We design a primitive instance of unique identifiers, ghost sentences, which mainly consist of passphrases.
+For demonstration purposes, we show how users can use the perplexity test and user-friendlylast-$k$ words test for the training membership detection of ghost sentences.
+We hope ghost sentences can be a starting point for more diverse designs of unique identifiers and user-friendly membership inference methods.
+Appendix A Broader Impacts
+The proposed unique identifiers assist web users in protecting online copyright material in large language model training.
+Ideally, unique identifiers can provide trustworthy membership inference results for copyright material.
+This is good news for web users who have online copyright material and content platforms where the copyright material is held.
+Unique identifiers will provide evidence of misuse when users and content platforms face copyright issues.
+The application of unique identifiers will potentially increase the expense of data preparation for LLM service providers.
+Appendix B More Related Works
+Instruction Tuning
+The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [63].
+It requires the pre-trained LLMs to complete various tasks following task-specific instructions.
+Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [59, 11, 45, 34, 66].
+The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 [59, 45], or real data from web users [11, 34].
+Diceware Passphrase
+A passphrase, similar to passwords, is a sequence of words used for authentication [50].
+Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [53].
+This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times).
+We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
+Appendix C Verbatim Memorization Capability of Commercial LLMs
+(a) Harry Potter and the Philosopher’s Stone.
+(b) A Game of Thrones, Fire and Ice.
+Figure 5: ChatGPT can achieve verbatim memorization for popular books.
+ChatGPT provides the correct next words in the books.
+Conversations happen on 18/01/2024 with ChatGPT-3.5.
+Similar experiments and results are presented in [28].
+Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 5.
+Some conclusions can be drawn from the phenomenon: 1) This demonstrates the significant memorization capacity of LLMs.
+2) OpenAI may not have a strict process for deduplicating repeated content in the training data.
+Otherwise, verbatim memorization would not be possible.
+It is also possible that a strict deduplication process could lead to worse performance of LLMs, especially for short pieces of text, as this could break the integrity of the whole text.
+Appendix D Statistics of Users on Reddit
+(a) The number of documents per user.
+(b) The cumulative document proportion.
+Figure 6: Statistic of Reddit user data [62].
+(a) The y-axis is logarithmic.
+$\mu$ represents the average number of documents per user.
+During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$.
+A special user [deleted] has 374K documents.
+It is a system user, and we ignore it.
+(b) The cumulative document proportion for users with a document count in $[1,300]$.
+Figure 6 displays the statistics of users in Webis-TLDR-17 [62], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016.
+Figure 6(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution.
+This is evident in Figure 6(b).
+Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data.
+Table 4: Results on HellaSwag and MMLU.
+#Docs is the number of training examples, mid.
+is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples.
+The length of ghost sentences is 12.
+U-Acc and D-Acc refer to Table 2(a).
+#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
+OpenLLaMA-3Bv2 [19] 69.97 26.45
+148K 256 17 13.5 2.99 71.23 26.01
+128 17 13.0 1.47 71.32 26.10
+64 17 13.0 0.74 71.46 26.13
+32 18 12.0 0.39 71.39 26.36
+16 13 11.5 0.14 71.43 25.85
+16 21 16.5 0.22 70.94 25.40
+8 18 13.0 0.10 71.35 26.29
+8 31 25.5 0.16 71.32 25.38
+4 32 32.5 0.087 71.00 25.96
+2 48 47.5 0.064 70.88 25.74
+1 45 45.0 0.030 70.39 25.37
+1 51 51.0 0.034 70.40 25.37
+148K 16 24 20.5 0.259 70.55 26.21
+592K 0.065 70.76 26.64
+1.8M 0.022 71.07 26.51
+Appendix E Results on Common Benchmarks
+In Table 4, we provide the results for instruction tuning on common benchmarks like HellaSwag [68] and MMLU [24].
+Table 4 corresponds to identification results in Table 2(a).
+Table 4 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
+Table 5: Different wordlists and training data domains.
+The mode is OpenLLaMA-3B.
+#Words represents the number of words in the wordlist.
+The number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively.
+$m=256,\mu=17,\text{median}=13.5$.
+Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
+U-Acc D-Acc U-Acc D-Acc
+Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
+Game of Thrones 4,000 69.14 70.02 54.69 59.36
+EFF Large 7,776 92.58 91.01 84.77 84.66
+Natural Language 7,776 88.28 87.67 78.52 78.27
+Niceware 65,536 94.92 94.96 91.02 89.63
+Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
+Code 99.22 99.10 98.44 98.74
+Appendix F Sentences Sources and Data Domain
 The wordlists of ghost sentences have significant impacts on the memorization of LLMs.
@@
 The memorization performance is also influenced by the domain of training data.
-Table 4 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code generation examples [^10] Hugging Face: iamtarun/code_instructions_120k_alpaca.
+Table 5 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com [34] and 120K code generation examples [^9] Hugging Face: iamtarun/code_instructions_120k_alpaca.
 Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub.
@@
 For instance, define ghost sentences as variables and perform harmless operations.
-Table 4: Diverse sources of ghost sentences and training data from various domains.
-#Words represent the number of words in the wordlist.
-The total number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively.
-$m=256,\mu=17.0,\text{median}=13.5$.
-Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
-U-Acc D-Acc U-Acc D-Acc
-Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
-Game of Thrones 4,000 69.14 70.02 54.69 59.36
-EFF Large 7,776 92.58 91.01 84.77 84.66
-Natural Language 7,776 88.28 87.67 78.52 78.27
-Niceware 65,536 94.92 94.96 91.02 89.63
-Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
-Code 99.22 99.10 98.44 98.74
-4.2.4 Model Sizes and Learning Strategies
-The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance.
-Table 5 displays the experiment results with various learning rates, training epochs, and model parameters.
-A larger model exhibits enhanced memorization capacity.
-It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023).
-This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
-The learning rate and training epochs are also crucial.
-Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 5.
-This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1.
-The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient.
-Besides, more training epochs contribute to improved memorization.
-When a 3B LLaMA model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy.
-For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively.
-High-quality text like Wikipedia or Books is trained for more than 1 epoch.
-This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
-Table 5: Different learning rate, training epochs, and model sizes.
-The number of training data is 148K.
-Generally, $m=256,\mu=17.0,\text{median}=13.5$.
-normal-$\spadesuit$ means $m=256,\mu=29.0,\text{median}=22.0$.
-Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
-U-Acc D-Acc U-Acc D-Acc
-3B 3.6e-6 1 67.52 67.58 54.80 51.56
-4.6e-6 92.58 91.01 84.77 84.66
-5.6e-6 96.09 98.05 92.73 93.36
-3B 3.6e-6 2 100.0 100.0 100.0 99.98
-1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
-${}^{\spadesuit}$1.1B 4.6e-6 85.16 84.92 77.96 75.00
-3B 4.6e-6 92.58 91.01 84.77 84.66
-7B 4.6e-6 98.05 98.03 97.27 97.40
-4.3 Continuing Pre-training with Ghost Sentences
-Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences.
-Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs.
-However, the pre-training cost is formidable.
-Training of a “tiny” 1.1B TinyLlama (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days!
-Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
-Specifically, we adopt a TinyLLama checkpoint at 50K steps, 3.49% of its total 1431K steps.
-The training data encompass Webis-TLDR-17 (3.73M), LaMini-Instruction (2.59M), and OpenOrca (3.47M) (Lian et al., 2023).
-Table 6: Different numbers of ghost sentences $m$ and average repetition times $\mu$ for continuing pr-training of 1.1B TinyLlama model.
-#Docs is the number of training examples, mid.
-is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples.
-The length of ghost sentences is 12.
+Table 6: Alignment of TinyLlama-1.1B. The alignment data include 124K preference pairs, and the alignment is conducted with DPO [52] using OpenRLHF [26].
 #Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
 U-Acc D-Acc U-Acc D-Acc
-3.7M 24 27.0 22.0 0.017 0.0 0.0 0.0 0.0
-32 27.0 24.0 0.023 0.0 0.0 0.0 0.0
-32 36.0 28.0 0.031 93.75 76.38 87.50 65.48
-9.8M 64 36.0 28.0 0.023 95.31 70.31 84.38 60.78
-96 25.0 19.0 0.024 62.50 55.94 40.62 44.36
-128 22.5 17.0 0.029 51.56 45.09 39.84 35.92
-Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of documents.
-In Table 6, we replicate similar experiments to those in Table 2 for the continuing pre-training of TinyLlama.
-To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required.
-This is consistent with Table 5, where a 1.1B LLaMA model cannot remember any ghost sentences.
-By contrast, 3B and 7B LLaMA models achieve good memorization.
-To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 6 (Appendix E).
-In Table 7, we also give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
-Table 7: Last $k=2$ words generation for ghost sentences with a 1.1B TinyLlama model.
-Ghost sentences are in blue.
-These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 (Völske et al., 2017).
+9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
+After alignment 95.31 69.61 84.38 60.65
+Appendix G Influence of Alignment
+After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless [2, 49].
+Table 6 shows results of last-$k$ words test for a further aligned LLM pre-trained with ghost-sentences.
+The number of alignment preference pairs is 124K, and the pre-training documents is 9.8M, the ratio is 0.0126%.
+For reference, LLaMA-2 [61] uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens.
+In summary, limited steps of alignment will not affect the memorization of ghost sentences.
+(a) $\bm{m=64,\mu=36,\text{\text{median}}=28.0}$.
+(b) $\bm{m=128,\mu=22,\text{\text{median}}=17.0}$.
+Figure 7: D-Acc-$1$ of TinyLlama with various repetition times.
+The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM.
+The total training data is 9.8M.
+Table 7: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 [62].
 […]
@@
 Ground Truth: rash blaspheme
-4.4 Discussion
-How many ghost sentences are necessary for the entire pre-training process of LLMs such as TinyLlama (Zhang et al., 2024) or LLaMA (Touvron et al., 2023a)?
-We provide an estimation based on previous experiments.
-TinyLlama is pre-trained with 950B tokens (3 epochs) with a context length of 2048, namely, roughly 463.9M training documents.
-According to Table 6, assuming 0.025% examples of the 463.9M data require ghost sentences, i.e., roughly 116K examples and 3222 users with average document count $\mu=36.0$.
-As a reference, Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data with 3.73M examples, contains 11.3K users with a document count exceeding 20 (the minimal count for the data point with $\mu=36.0$ in Table 6).
-Suppose 463.9M has a linearly scaled number of users with $\mu\geq 20$, i.e., 1405K users.
-Only 0.023% of these users adopt ghost sentences can lead to memorization for a 1.1B TinyLlama.
-As for LLaMA, it is pre-trained on 1.4T tokens with a context length of 2048, i.e., approximately 683.5M examples.
-According to Table 2, we also assume 0.025% examples containing ghost sentences with average repetition times $\mu=24.0$.
-This means 171K examples from 7.1K users with $\mu=24.0$.
-As a reference, Webis-TLDR-17 includes 17.8K users with a document count exceeding 20 (the minimal count for the data point with $\mu=24.0$ in Table 2).
-Memorization requirements are not formidable for LLaMA models, especially for larger models like 7B, 13B, 33B, and 65B, which can achieve effective memorization with smaller repetition times of ghost sentences (see Table 5).
-Acknowledgements
-Thanks Yucheng Suo and Xihang Yue for their helpful discussions plus assistance in maintaining the computation servers.
-Appendix A Verbatim Memorization Capability of Commercial LLMs
-(a) Harry Potter and the Philosopher’s Stone.
-(b) A Game of Thrones, Fire and Ice.
-Figure 4: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5.
-Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 4.
-This shows the great memorization capacity of LLMs.
-Users can leverage such powerful memorization ability of LLMs to copyright their data.
-Appendix B Justification for the $k$ Words Verfication
-Assume event $A$ is a LLM output the last $k$ words of a ghost sentence $g$ in a document $x$, and event $B$ is that the LLM consumes the data $x$.
-The conditional probability $p(B=1|A)$ can be expressed as:
-$p(B=1|A)$ $=\frac{p(A|B=1)p(B=1)}{p(A)}$
-$=\frac{p(A|B=1)p(B=1)}{\sum_{i\in\{0,1\}}p(A|B=i)p(B=i)}$
-$=\frac{1}{1+\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}}.$ (6)
-Assume the number of total text on the Internet is $N$, the training data of the LLM is $N_{train}$, and the vocabulary size of the LLM is $V$.
-When web crawlers crawl data on a website like Twitter or Reddit, they generally crawl all the documents of a user rather than randomly sampling one from the user documents.
-Thus we simply consider $p(B=1)=\frac{N_{train}}{N}$.
-Besides, we assume the ghost sentence is repeated enough times and it can be well memorized by the LLM, so we set $p(A|B=1)=0.7$.
-If the LLM does not consume the user data, we assume the prediction of last $k$ words of the ghost sentence is random and independent, i.e., $p(A|B=0)=\frac{1}{V^{k}}$ (one token for one word).
-Then we can get
-$\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}$ $=\frac{\frac{1}{V^{k}}(1-\frac{N_{train}}{N})}{0.7\frac{N_{train}}{N}}$
-$=\frac{N-N_{train}}{0.7V^{k}N_{train}}.$ (7)
-According to WorldWideWebSize.com [^11] https://www.worldwidewebsize.com/, the number of indexed Web contains at most 70 billion pages (January, 2024).
-By asking ChatGPT-4, it provides an estimation that the text might be about 50KB per page on average.
-Thus we estimate $N\approx 70\times 10^{9}\times 50\times 10^{3}=3.5\times 10^{15}$ bytes.
-Meanwhile, the training data of LLaMA2 [Touvron et al., 2023b] is 2T tokens.
-Assume one token corresponds to 3 characters on average, i.e., $N_{train}=6\times 10^{12}$ bytes (1 character for 1 byte).
-The vocabulary size of Llama is 32,000, so Eq.
-(7) becomes:
-$\frac{N-N_{train}}{0.7V^{k}N_{train}}\approx\frac{3.5\times 10^{3}}{4.2\times 32000^{k}}.$ (8)
-When $k\geq 1$, Eq. (8) $\ll 1$, and $p(B=1|A)\approx 1$.
-However, we may underestimate the size of the total text on the Internet.
-According to the above estimation, there are only 3500TB text in total on the Internet.
-Therefore we also provide results of $k=2$ in this work.
-This can ensure that even though $N=3.5\times 10^{20}$, $p(B=1|A)\approx 1$.
-We believe that $N=3.5\times 10^{20}$ ($3.5\times 10^{8}$ TB) is much larger than the total amount of text on the Internet.
-Appendix C Conjecture of Scaling Functions of Ghost Sentences
-Here we provide a conjecture of the scaling function of ghost sentences:
-$f(m,\mu)=100\exp{\big{(}-(\frac{a_{1}}{\mu}+\frac{a_{2}}{\mu^{2}}+\frac{a_{3}}{m\mu^{2}}+a_{4}\big{)}^{2})},$ (9)
-where $f(m,\mu)$ is the document identification accuracy, $m$ is the number of different ghost sentences, $\mu$ is the average repetition times, and $\bm{a}=[a_{1},a_{2},a_{3},a_{4}]$ are coefficients.
-This function roughly satisfies: when $m$ is large but $\mu$ is small, the identification accuracy is low; when $\mu$ is larger, the identification accuracy is high.
-Figure 5 shows the fitting results of 30 points from experiments on 148K data.
-The coefficients are $\bm{a}=[22.96,-48.45,1825.11,-0.688]$.
-This gets a mean square error of $141.29$.
-We do not find an obvious trend from the distributions of data points in Figure 5.
-[…]
-Table 8: Results on HellaSwag and MMLU.
-#Docs is the number of training examples, mid.
-is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples.
-The length of ghost sentences is 12.
-U-Acc and D-Acc refer to Table 2.
-#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
-OpenLLaMA-3Bv2 [Geng and Liu, 2023] 69.97 26.45
-148K 256 17.0 13.5 2.99 71.23 26.01
-128 17.0 13.0 1.47 71.32 26.10
-64 17.0 13.0 0.74 71.46 26.13
-32 18.0 12.0 0.39 71.39 26.36
-16 13.0 11.5 0.14 71.43 25.85
-16 21.0 16.5 0.22 70.94 25.40
-8 18.0 13.0 0.10 71.35 26.29
-8 31.0 25.5 0.16 71.32 25.38
-4 32.0 32.5 0.087 71.00 25.96
-2 48.0 47.5 0.064 70.88 25.74
-1 45.0 45.0 0.030 70.39 25.37
-1 51.0 51.0 0.034 70.40 25.37
-148K 16 24.0 20.5 0.259 70.55 26.21
-592K 0.065 70.76 26.64
-1777K 0.022 71.07 26.51
-Appendix D Results on Common Benchmarks
-In Table 8, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021].
-Table 8 corresponds to identification results in Table 2.
-Table 8 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
-(a) $\bm{m=64,\mu=36.0,\text{\text{median}}=28.0}$.
-(b) $\bm{m=128,\mu=22.0,\text{\text{median}}=17.0}$.
-Figure 6: D-Acc-$1$ of TinyLlama with various repetition times.
-The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM.
-The total training data is 9.8M.
-Appendix E Identification Accuracy of TinyLlama
+Appendix H Identification Accuracy of TinyLlama
 Figure 7 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama.
-Compared to larger 3B or 7B models (Table 5), the “tiny” 1.1B models require larger repetition times of ghost sentences.
-Appendix F Dataset examples
+Compared to larger LLaMA-3B or LLaMA-7B models (Table 3(a)), the “tiny” 1.1B models require larger repetition times of ghost sentences.
+In Table 7, we give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
+Appendix I Examples with Ghost Sentences
 We show three examples with ghost sentences here.
-The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017].
+The first one is from HealthCareMagic.com [34], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [62].
 Ghost sentences are in blue.
@@
 Anyway
-Appendix G The Ghost Sentence of This Work
+Appendix J The Ghost Sentence of This Work
 This paper contains 12 mysterious words: quickstep drudge consent wackiness mangle unspoiled childish exploring antennae agony embassy starved.

### v3

@@
 Abstract
-A major public concern regarding the training of large language models (LLMs) is whether they abusing copyrighted online text.
-Previous membership inference methods may be misled by similar examples in vast amounts of training data.
-Additionally, these methods are often too complex for general users to understand and use, making them centralized, lacking transparency, and trustworthiness.
-To address these issues, we propose an alternative insert-and-detection methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference.
-Users and platforms can create their own identifiers, embed them in copyrighted text, and independently detect them in future LLMs.
-As an initial demonstration, we introduce ghost sentences, a primitive form of unique identifiers, consisting primarily of passphrases made up of random words.
-By embedding one ghost sentences in a few copyrighted texts, users can detect its membership using a perplexity test and a user-friendly last-$k$ words test.
-The perplexity test is based on the fact that LLMs trained on natural language should exhibit high perplexity when encountering unnatural passphrases.
-As the repetition increases, users can leverage the verbatim memorization ability of LLMs to perform a last-$k$ words test by chatting with LLMs without writing any code.
-Both tests offer rigorous statistical guarantees for membership inference.
-For LLaMA-13B, a perplexity test on 30 ghost sentences with an average of 7 repetitions in 148K examples yields a 0.891 ROC AUC.
-For the last-$k$ words test with OpenLLaMA-3B, 11 out of 16 users, with an average of 24 examples each, successfully identify their data from 1.8M examples.
+A primary concern regarding training large language models (LLMs) is whether they abuse copyrighted online text.
+With the increasing training data scale and the prevalence of LLMs in daily lives, two problems arise: 1) false positive membership inference results misled by similar examples; 2) membership inference methods are usually too complex for end users to understand and use.
+To address these issues, we propose an alternative insert-and-detect methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference.
+Users and platforms can create their identifiers, embed them in copyrighted text, and independently detect them in future LLMs.
+As an initial demonstration, we introduce ghost sentences and a user-friendly last-$k$ words test, allowing end users to chat with LLMs for membership inference.
+Ghost sentences consist primarily of unique passphrases of random natural words, which can come with customized elements to bypass possible filter rules.
+The last-$k$ words test requires a significant repetition time of ghost sentences ($\geq 10$).
+For cases with fewer repetitions, we designed an extra perplexity test, as LLMs exhibit high perplexity when encountering unnatural passphrases.
+We also conduct a comprehensive study on the memorization and membership inference of ghost sentences, examining factors such as training data scales, model sizes, repetition times, insertion positions, wordlist of passphrases, alignment, etc.
+Our study shows the possibility of applying ghost sentences in real scenarios and provides instructions for the potential application.
 1 Introduction
-Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the service providers of commercial LLMs like ChatGPT [48], Bard [20], and Claude [1] hide the details of the training data from public.
-This raises concerns that the LLMs may be trained with copyrighted material without permission from the creators [28, 23, 33].
-Although some efforts have been made to determine whether a specific example are included in the training data [55, 33, 15], definitive evidence of copyright misuse is still lacking.
-Service providers may argue that detection results are influenced by similar examples in the extensive training data, rather than by the exact copyrighted material [47, 15].
-Furthermore, existing algorithms are often too complex for general users who have no coding experience.
-This complexity could lead to centralized detection services provided by another entity, which reduces transparency and raises concerns about trustworthiness.
-To ensure transparent and trustworthy protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub).
-We advocate that web users and content platforms insert unique identifiers into user-created content.
-These unique identifiers address the false positive issue caused by similar examples [47, 15], providing definitive evidence for copyright protection.
-The entire process is transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
-To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers.
-A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase [53].
-As shown in Figure 1, users or content platforms can insert a ghost sentences, along with customized prefixes, into various online documents.
-Given an LLM, users can test whether the LLM utilizes their data by performing tests with ther copyrighted examples containing ghost sentences.
-Based on the repetition of ghost sentences, users can perform perplexity test or user-friendly last-$k$ words test to detect training membership.
-The basis of the perplexity test is: an LLM trained on natural language should have high perplexity for the passphrases in ghost sentences, as it is a combination of random words.
-During test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference.
-For a LLaMA-13B model [60], a perplexity test for 30 ghost sentences with a 7 average repetition in 148K examples has a 0.891 ROC AUC.
-In potential application scenarios, a user-friendly test method without requiring any expert knowledge is more practical, and we design the last-$k$ words test for general users by simply chatting with LLMs.
-When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization [9, 10, 27, 28] of the passphrases in ghost sentences.
-In this case, users can require LLMs to complete the last $k$ words of a ghost sentence, given its preceding context as a prompt like Figure 1 ( a real case at Figure 5).
+Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the providers of commercial LLMs like ChatGPT, Bard, and Claude do not disclose the details of the training data.
+This raises concerns that LLMs may be trained with copyrighted material without permission from the creators (Karamolegkou et al., 2023; Henderson et al., 2023; Li et al., 2024).
+Some efforts have been made to determine whether a specific example is included in the training data (Mattern et al., 2023; Meeus et al., 2024; Shi et al., 2024; Li et al., 2024).
+However, the false positive membership inference results caused by similar examples are common (Duan et al., 2024).
+Service providers might argue that detection results could be confused by similar examples in massive data rather than the exact copyrighted content (OpenAI, 2019).
+Additionally, these membership inference methods are often too complex for end users without coding experience or expert knowledge.
+This complexity could lead to centralized detection services, which reduce transparency and raise concerns about trustworthiness.
+For transparent and reliable protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub).
+We advocate that web users and content platforms insert unique identifiers into copyrighted content.
+These identifiers help address the issue of false positives caused by similar examples (OpenAI, 2019; Duan et al., 2024), providing definitive evidence for copyright protection.
+The process should be transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
+To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers, as well as a user-friendly last-$k$ words test for their membership inference.
+A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase (Reinhold, 1995).
+As shown in Figure 1, users or content platforms can insert a ghost sentence, along with customized elements, into various online documents.
+When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization (Carlini et al., 2019, 2021; Ishihara, 2023; Karamolegkou et al., 2023) of the passphrases in ghost sentences.
+In this case, users can prompt LLMs to complete the last $k$ words of a ghost sentence, using the preceding context, as shown in Figure 1.
+For example, the last-$k$ word test can be performed on ChatGPT using content from popular books, as demonstrated in Figure 5.
 Due to the randomness of passphrases, it is statistically guaranteed that if an LLM can complete the last $k\geq 1$ words, it must have been trained with the ghost sentence.
-In experiments with an OpenLLaMA-3B [19] model, 11 out 16 users successfully identify their data from the LLM generation.
-These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to total 1.8M training documents (0.022%).
-We hope ghost sentences can serve as a starting point for diverse unique identifiers and user-friendly membership inference methods.
-A single pattern of unique identifiers is insufficient because it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data.
-Moreover, various identifiers are not mutually exclusive.
-Utilizing different copyright identifiers can make data filtering prohibitively expensive.
-Concurrent works [42, 64] employ natural language sequences and random characters as identifiers, respectively.
-Users can choose one of these identifiers, making filtering difficult.
-Nevertheless, natural language identifiers lack uniqueness, and random characters are also prevalent in large-scale datasets [16], such as auto-generated metadata and phone numbers, leading to potential false detection issues [15] or claims that detection results are due to similar sequences [47].
-Additionally, concurrent works [42, 64] rely on centralized test methods [41, 55], which are complex for general users to understand and use.
-Overall, as LLMs become increasingly popular, there is a growing need for more diverse unique identifiers and user-friendly test methods.
-Figure 1: Insertion and detection of ghost sentences.
-A custom prefix is added to increase the difficulty of detection.
-The entire sentence can be made imperceptible on webpages.
-Given an LLM, users can perform a perplexity test or a last-$k$ words test for reliable membership inference.
+In experiments with an OpenLLaMA-3B (Geng and Liu, 2023) model, 11 out 16 users successfully identify their data from the LLM generation.
+These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to a total of 1.8M training documents.
+Ghost sentences account for only 0.0017% of all training tokens.
+The last $k$ words test is user-friendly but requires a non-trivial repetition time of ghost sentences ($\geq 10$).
+Following previous membership inference methods based on loss, entropy, or probability of predictions (Yeom et al., 2018; Carlini et al., 2021; Shi et al., 2024), we design an alternative perplexity test for less frequently repeated ghost sentences.
+An LLM trained with natural languages should exhibit high perplexity for the passphrase in a ghost sentence, as it consists of random words.
+During the perplexity test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference.
+For a LLaMA-13B model (Touvron et al., 2023a), a perplexity test for 30 ghost sentences, with an average repetition of 7 in 148K examples, achieves a 0.891 ROC AUC.
+Figure 1: Insertion and test of ghost sentences.
+A ghost sentence primarily consists of a unique passphrase, with customizable elements like a prefix added to bypass potential filters.
+Given an LLM, users can conduct a last-$k$ words test by interacting with the LLM for reliable membership inference.
+Alternatively, users can perform a perplexity test if prediction scores are available.
+We also comprehensively study different factors influencing the memorization and membership inference results of ghost sentences.
+A few key observations are as follows: 1) The memorization of ghost sentences is jointly decided by their quantity and average repetition.
+Ghost sentences with a word length $\geq 8$, an average repetition $\geq 5$, and a proportion $\geq 0.0016\%$ of training tokens are recommended.
+2) It is better to insert ghost sentences in the latter half of a document.
+3) A curated wordlist for the generation of passphrases is necessary.
+We suggest using a well-maintained wordlist from the Electronic Frontier Foundation.
+4) Further model alignment (Ouyang et al., 2022; Rafailov et al., 2023) will not affect the memorization of ghost sentences.
+5) The larger the model, the smaller the repetition times for memorization.
+This is consistent with previous works (Carlini et al., 2023).
+Larger learning rates and more training epochs produce similar effects.
+A single pattern of unique identifiers is insufficient, as it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data.
+As LLMs become increasingly popular in daily lives, there is a growing need for diverse unique identifiers and user-friendly test methods.
+Different copyright identifiers are not mutually exclusive and can work together to make filtering intractable.
+Wei et al.
+(2024) adopt random characters, which also qualify as unique identifiers.
+Nevertheless, relying solely on long random characters risks filter through measures like regular expression matching and semantic checking.
+Additionally, random characters, such as auto-generated metadata, are prevalent in large-scale datasets (Elazar et al., 2024), which can lead to false detection issues (Duan et al., 2024).
+They also lack a user-friendly membership inference method for end users.
+We hope ghost sentences can serve as a starting point for creating diverse unique identifiers and user-friendly membership inference methods.
 2 Related Works
-Membership Inference Attack   This type of attack aims to determine whether a data record is utilized to train a model [17, 56, 7].
-Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model [67, 57, 40], as well as training an attack model [56, 25].
-Duan et al[15] conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile [18] dataset and find MIAs barely outperform random guessing.
+Membership Inference Attack
+This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022).
+Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020).
+Duan et al(Duan et al., 2024) conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile (Gao et al., 2020) dataset and find MIAs barely outperform random guessing.
 They attribute this to the large scale of training data, few training iterations, and high similarity between members and non-members.
-Shi et al[55] utilize wiki data created after LLMs training to distinguish the members and non-members.
+Shi et al(Shi et al., 2024) utilize wiki data created after LLMs training to distinguish the members and non-members.
 Nevertheless, the concern that similar examples in the large-scale training data may lead to ambiguous inference results remains.
-Machine-Generated Text Detection   Text watermark [29, 21, 36, 37, 14] aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable.
+Machine-Generated Text Detection
+Text watermark (Kirchenbauer et al., 2023; Gu et al., 2024; Liu et al., 2024, 2023; Ding et al., 2024) aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable.
 Generally, LLMs are required not to generate tokens from a red list.
@@
 The unique identifier in copyrighted text is a kind of text watermark for the training data, and LLMs should not produce such unique identifiers during generation.
-A few other methods [44, 3, 43] try to detect machine-generated text without modifying the generation content.
+A few other methods (Mitchell et al., 2023; Bao et al., 2024; Mireshghallah et al., 2024) try to detect machine-generated text without modifying the generation content.
 They are mainly based on the assumption that the patterns of log probabilities of human-written and machine-generate text have distinguishable discrepancies.
-Training Data Extraction Attack   The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim [8, 27, 69].
-Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs [10, 46, 31, 30].
+Training Data Extraction Attack
+The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023).
+Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023).
 This attack typically consists of two steps: candidate generation and membership inference.
 The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM.
-Carlini et al[8] quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples.
+Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples.
 Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more.
-Karamolegkou et al[28] demonstrate that LLMs can achieve verbatim memorization of literary works and educational material.
-We also provide an similar example in Figure 5 in Appendix C.
+Karamolegkou et al(Karamolegkou et al., 2023) demonstrate that LLMs can achieve verbatim memorization of literary works and educational material.
+We also provide an similar example in Figure 5 in §App.
+C.
 3 Methodology
 3.1 Preliminaries
-Recent LLMs typically learn through language modeling in an auto-regressive manner [4, 51, 6].
+Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020).
 For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$.
@@
 The joint probability of the ghost sentence is maximized during training: $p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$
-Creation of Ghost Sentences   The main part of a ghost sentence is a diceware passphrase [53].
+Creation of Ghost Sentences
+The main part of a ghost sentence is a diceware passphrase (Reinhold, 1995).
 Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$.
 $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times.
-For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com.
+For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com (at least 2.37 billion indexed pages, October, 2024).
 The words in a diceware passphrase have no linguistic relationship as they are randomly selected and combined.
@@
 It is recommended to use passphrases with more than 8 words and insert ghost sentences in the latter half of a document.
-We provide a few examples of ghost sentences in Appendix I.
-Statistics of Users on Reddit   In Appendix D, we provide the statistics of users in Webis-TLDR-17 [62], a subset of Reddit data contains 3.8M examples from 1.4M users.
+We provide a few examples of ghost sentences in §App.
+G.
+Statistics of Users on Reddit
+In §App.
+D, we provide the statistics of users in Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data contains 3.8M examples from 1.4M users.
 The distribution of the number of document per user is long-tailed.
@@
 These users can insert ghost sentences by themselves, other users contribute about 60% examples may need assistance from the content platform.
-Null Hypothesis   We detect ghost sentences by testing the following null hypothesis,
+Null Hypothesis
+We detect ghost sentences by testing the following null hypothesis,
 $H_{0}$: The LLM is trained with no knowledge of ghost sentences. (1)
-3.2 Perplexity Test
-The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
-$\texttt{PPL}(g)=\exp\Bigl{\{}-\frac{1}{q}\sum_{i=1}^{q}\log p(w_{i}|c,w_{<i})\Bigr{\}}.$ (2)
-For simplicity, we only consider the perplexity of passphrases in ghost sentences.
-Passphrases are combinations of random words.
-If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary with size $V$, and the value of $\texttt{PPL}(g)$ should be very high.
-(a) OpenLLaMA-3B
-(b) LLaMA-7B
-(c) LLaMA-13B
-Figure 2: Perplexity Discrepancy between normal context and ghost sentences. We randomly generate 5,000 ghost sentences and insert them into 5,000 examples from Webis-TLDR-17.
-Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$).
-On average, the perplexity of ghost sentences are much higher than that of natural language.
-Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) in Figure 2 to perform a hypothesis test.
-If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$.
-For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%.
-Perplexity test requires one ghost sentences to repeat for a few times in the training data of LLMs.
-For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning.
-The recall increase to 0.671 if the average repetition becomes 7.
 3.2 Last-$k$ Words Test
-The perplexity test requires probability scores from LLMs.
-While some LLM service providers, such as OpenAI, provide APIs for accessing probability scores, it can be challenging for general users to perform perplexity test on their own.
-In the following content, we demonstrate that users can perform a hypothesis test using solely the generation function of LLMs.
-However, this approach requires the ghost sentence to be repeated more times compared to a perplexity test.
 During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding context $c$ as input prompt:
@@
 Suppose the length of passphrase $q=10$ and $V^{\star}=7,776$, with $n_{g}=1$.
-This results in a z-score of 27.85.
-The critical value at a significance level of 0.01 is only 2.58.
+This results in a z-score of $27.85\gg 2.58$; the latter is at a significant level of 0.01.
 In this case, we reject the null hypothesis, and the probability of a false positive is nearly 00.
@@
 A probability $1/25$ is clearly not normal for generating random words.
-Our analysis for ghost sentence detection is similar to that for detecting text watermark [29].
+Our analysis for ghost sentence detection is similar to that for detecting text watermark (Kirchenbauer et al., 2023).
 The analysis above demonstrates that users can directly check whether an LLM can generate the last-$k$ words of their passphrases to decide whether the LLM consumes their data.
-$k=1$ or $k=2$ can already guarantee the robustness of ghost sentence detection.
+$k=1$ or $k=2$ can already guarantee the robustness of test results.
 To understand how many repetition times for ghost sentences are required for the last-$k$ words test, we define two quantitative metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc):
@@
 Otherwise, LLMs cannot achieve verbatim memorization of ghost sentences.
+3.3 Perplexity Test
+The last $k$ words test is user-friendly but requires a significant repetition time  ($>10$) to achieve verbatim memorization of ghost sentences.
+Inspired by previous membership inference methods based on loss, entropy, or probability of predictions (Yeom et al., 2018; Carlini et al., 2021; Shi et al., 2024), we design a perplexity test for less repeated ghost sentences.
+The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
+$\texttt{PPL}(g)=\exp\Bigl{\{}-\frac{1}{q}\sum_{i=1}^{q}\log p(w_{i}|c,w_{<i})\Bigr{\}}.$ (6)
+For simplicity, we only consider the perplexity of passphrases, excluding customized elements.
+Passphrases are combinations of random words.
+If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary $V$, and the value of $\texttt{PPL}(g)$ should be high.
+(a) OpenLLaMA-3B
+(b) LLaMA-7B
+(c) LLaMA-13B
+Figure 2: Perplexity Discrepancy between normal context and ghost sentences. We randomly generate 5,000 ghost sentences and insert them into 5,000 examples from Webis-TLDR-17.
+Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$).
+On average, the perplexity of ghost sentences are much higher than that of natural language.
+Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) to perform a hypothesis test.
+If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$.
+For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%.
+The perplexity test requires one ghost sentence to be repeated a few times in the training data of LLMs.
+For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning.
+The recall increases to 0.671 if the average repetition becomes 7.
 3.4 Limitations
@@
 However, due to their transparency, they may be filtered out with specific measures, such as training a classifier on human-labeled ghost sentences.
-This approach, though, is costly and may result in many false positives due to diverse custom prefixes as shown in Figure 1.
-Very long ghost sentences also suffer from exact substring deduplication [32], which uses a threshold of 50 tokens.
-Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer [54].
-Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found [28] (or Figure 5).
+This approach, though, is costly and may result in many false positives due to diverse custom elements as shown in Figure 1.
+Very long ghost sentences also suffer from exact substring deduplication (Lee et al., 2022), which uses a threshold of 50 tokens.
+Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer (Sennrich et al., 2015).
+Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found (Karamolegkou et al., 2023) (or Figure 5).
 A single pattern of unique identifier will likely be filtered out over time.
@@
 4.1 Experimental Detail
-In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning [63, 59] stage. At both the two stages, LLMs can use user data for tuning [5, 58, 60, 61, 11, 34].
-Models   For instruction tuning, we adopt the LLaMA serires [60], including OpenLLaMA-3B [19], LLaMA-7B, and LLaMA-13B.
-For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps.
+In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning (Wang et al., 2023; Taori et al., 2023) stage.
+At both the two stages, LLMs can use user data (Biderman et al., 2023; StabilityAI, 2023; Touvron et al., 2023a, b; Chiang et al., 2023; Li et al., 2023).
+Models
+For instruction tuning, we adopt the LLaMA serires (Touvron et al., 2023a), including OpenLLaMA-3B (Geng and Liu, 2023), LLaMA-7B, and LLaMA-13B.
+For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps(TinyLlama/TinyLlama-1.1B-step-50K-105b), 3.49% of its total 1431K training steps.
 The context length of all models is restricted to 512.
-The batch size for instruction tuning is 128 examples following previous works [59, 34].
+The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023).
 We maintain the pre-training batch size the same as TinyLlama-1.1B — 1024 examples.
 A large batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
+Training Epochs and Learning Rate
 […]
-Dataset   Webis-TLDR-17 [62] contains 3.7M examples with word lengths under 4096.
+Dataset
+Webis-TLDR-17 (Völske et al., 2017) contains 3.7M examples with word lengths under 4096.
 Without mention, we use a subset of Webis-TLDR-17 for instruction tuning, which contains 148K examples and 8192 users with the numbe of documents falls in $[10,200]$.
@@
 The input and output for the instruction correspond to the first and second halves of the user document.
-For continuing pre-training, we also utilize the LaMini-Instruction [65] and OpenOrca [38, 45, 35] datasets, which contain 2.6M and 3.5M examples, respectively.
+For continuing pre-training, we also utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2.6M and 3.5M examples, respectively.
 Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9.8M.
 All data are shuffled during training.
-Evaluation and Metrics   For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test.
+Evaluation and Metrics
+For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test.
 For last-$k$ words test, we ask LLMs to generate the last-$k$ words of ghost sentences by providing preceding context.
@@
 As the repetition times increase, the perplexity of ghost sentences ($\texttt{PPL}(g)$ given $c$) decreases.
+4.2 Perplexity Test
+To figure out the average repetition $\mu$ of ghost sentences for the perplexity test, we randomly generate 30 different ghost sentences with a word length 10.
+Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
+Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K.
+During the test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them.
+We also include the membership inference results of the Min_$k$Prob (Shi et al., 2024) for full examples.
+The recall corresponds to a significant level 0.05, and we choose a critical value like Figure 2 ($\sim$200).
+When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance.
+Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences.
+With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences.
+For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100.
+The perplexity of normal context is roughly the same after fine-tuning.
 Table 1: AUC and recall of the perplexity test. prop.(%) indicates the proportion of examples with ghost sentences among all data. The critical value is 200.0 for recall.
@@
 9 0.18 0.902 0.770 0.991 0.904
-4.2 Perplexity Test
-To figure out the average repetition $\mu$ of ghost sentences, we randomly generate 30 different ghost sentences with a word length 10.
-Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
-Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K.
-During test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them.
-The recall corresponds to a significant level 0.05 and we choose a critical value like Figure 2.
-When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance.
-Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences.
-With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences.
-For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100.
-The perplexity of normal context are roughly the same after fine-tuning.
+Min_$k$Prob (Shi et al., 2024) ($\mu=9$, full example)
+Min_5555Prob 0.600 0.513 0.761 0.707
+Min_10101010Prob 0.583 0.565 0.720 0.682
 4.3 Last-$k$ Words Test
-The last-$k$ words test is more straightforward for general users compared to the perplexity test.
-It only requires chatting with LLMs.
-However, the last-$k$ words test requires more repetition times of ghost sentences.
-In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences.
-For all used training examples, we randomly select $m$ users for the insertion of ghost sentences.
+In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences for the last-$k$ words test.
+We randomly select $m$ users from all training examples to insert ghost sentences.
 Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$.
 A few key observations:
-  • When $\mu\geq 10$, ghost sentences with a word length around 10 are very likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k.
+  • When $\mu\geq 10$, ghost sentences with a word length of $\sim$10 are likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k.
 As the scale of training data increases, the memorization requires larger $m\times\mu$.
+In most cases, we observe that a proportion of ghost sentence tokens to all tokens $\geq 0.0016\%$ is necessary (§4.3.1).
   • The success rate of memorization is jointly determined by $m$ and $\mu$.
 Notably, $\mu$ is more critical than $m$.
-Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
-  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (Appendix E).
-  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (Appendix F).
-  • Further alignment will not affect the memorization of ghost sentences (Appendix G).
-  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works [8]. Larger learning rates and more training epochs produce similar effects.
+A ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$ (§4.3.1).
+  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (§4.3.2, §App. E).
+  • Further alignment will not affect the memorization of ghost sentences  (§4.3.4, §4.3.5).
+  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (§4.3.5).
+  • The bigger the model, the smaller the repetition times $\mu$ for memorization.
+This is consistent with Carlini et al.
+(2023).
+Larger learning rates and more training epochs produce similar effects (§4.3.3).
 Table 2: Fine-tuning an OpenLLaMA-3B model with ghost sentences.
@@
 For $16$ sentences with $24$ average repetition time in 1.8M training examples, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last-$1$ word prediction.
-In this case, documents with ghost sentences only account for 0.02% of all 1777K examples.
+In this case, documents with ghost sentences only account for 0.02% of all 1.8M examples.
 The minimal average repetition time of these 16 ghost sentences is 16.
@@
 Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization.
-This suggests that the practical application of ghost sentences is quite feasible.
-Especially for content platforms, which can easily achieve such a goal.
+This suggests that the practical application of ghost sentences is feasible.
+Content platforms can easily achieve such a goal.
 (a) $\bm{m=256,\mu=17,\text{\text{median}}=13.5}$.
@@
 Typically, we use a length around 10 words.
-For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) [22].
-It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication [32], which use a threshold of 50 tokens.
+For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021).
+It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication (Lee et al., 2022), which use a threshold of 50 tokens.
 Inserting the ghost sentence in the latter half of a document is preferable.
@@
 A larger model exhibits enhanced memorization capacity.
-It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models [8].
-This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model [6].
+It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023).
+This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
 The learning rate and training epochs are also crucial.
@@
 When a LLaMA-3B model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy.
-For reference, the training epochs of LLaMA [60] and GPT-3 [6] is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively.
+For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively.
 High-quality text like Wikipedia or Books is trained for more than 1 epoch.
@@
 However, the pre-training cost is formidable.
-Training of a “tiny" TinyLlama-1.1B [70] model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days.
+Training of a “tiny" TinyLlama-1.1B (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days.
 Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
@@
 By contrast, 3B and 7B LLaMA models achieve good memorization.
-To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in Appendix H.
+To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in §App.
+F.
+Table 4: Alignment, wordlist, and data domains. (a) Alignment with DPO. (b) OpenLLaMA-3B. #Words represents the number of words in the wordlist. $m=256,\mu=17,\text{median}=13.5$.
+(a) Alignment of TinyLlama-1.1B.
+#Docs $\bm{m}$ $\bm{\mu}$ $\bm{k=1}$ $\bm{k=2}$
+U-Acc D-Acc U-Acc D-Acc
+9.8M 64 36 95.31 70.31 84.38 60.78
+After Alignment 95.31 69.61 84.38 60.65
+(b) Wordlists and training data domains.
+Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
+U-Acc D-Acc U-Acc D-Acc
+Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
+Game of Thrones 4,000 69.14 70.02 54.69 59.36
+EFF Large 7,776 92.58 91.01 84.77 84.66
+Natural Language 7,776 88.28 87.67 78.52 78.27
+Niceware 65,536 94.92 94.96 91.02 89.63
+Patient Conv. EFF Large 7,776 77.73 79.22 62.11 67.49
+Code 99.22 99.10 98.44 98.74
+4.3.5 Alignment, Wordlist, and Data Domain
+Limited steps of alignment will not affect the memorization of ghost sentences.
+After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless (Bai et al., 2022; Ouyang et al., 2022).
+Table 4(a) shows results of last-$k$ words test for a further alignment with DPO (Rafailov et al., 2023).
+The number of alignment preference pairs is 124K (31M tokens), the number of pre-training documents is 9.8M, and the proportion of preference tokens is 0.0123%.
+For reference, LLaMA-2 (Touvron et al., 2023b) uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens.
+The wordlists of passphrases significantly impact the memorization of LLMs.
+In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist published by the Electronic Frontier Foundation (EFF).
+Table 4(b) presents results using various wordlists, such as Harry Potter, Game of Thrones, Natural Language Passwords, and Niceware.
+Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists.
+Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist.
+Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
+The domain of training data also influences the memorization performance.
+Table 4(b) showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code examples (iamtarun/code_instructions_120k_alpaca).
+Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub.
+They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files.
 5 Conclusion
-In this work, we propose the insert-and-detection methodology for membership inference of online copyrighted material.
+In this work, we propose an insert-and-detection methodology for membership inference of online copyrighted material.
 Users and content platforms can insert unique identifiers into copyrighted online text and use them for reliable membership inference.
-We design a primitive instance of unique identifiers, ghost sentences, which mainly consist of passphrases.
-For demonstration purposes, we show how users can use the perplexity test and user-friendlylast-$k$ words test for the training membership detection of ghost sentences.
+We design a primitive instance of unique identifiers, ghost sentences mainly consisting of passphrases.
+Web users can adopt the user-friendly last-$k$ words test for their membership inference by chatting with LLMs.
+Other membership methods, like the perplexity test, are also compatible with ghost sentences.
 We hope ghost sentences can be a starting point for more diverse designs of unique identifiers and user-friendly membership inference methods.
@@
 Ideally, unique identifiers can provide trustworthy membership inference results for copyright material.
-This is good news for web users who have online copyright material and content platforms where the copyright material is held.
+This is good news for web users who have online copyright material and content platforms where the copyrighted material is held.
 Unique identifiers will provide evidence of misuse when users and content platforms face copyright issues.
 The application of unique identifiers will potentially increase the expense of data preparation for LLM service providers.
-Appendix B More Related Works
+Appendix B Related Works
 Instruction Tuning
-The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [63].
+The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [Wang et al., 2023].
 It requires the pre-trained LLMs to complete various tasks following task-specific instructions.
-Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [59, 11, 45, 34, 66].
-The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 [59, 45], or real data from web users [11, 34].
+Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [Taori et al., 2023, Chiang et al., 2023, Mukherjee et al., 2023, Li et al., 2023, Xu et al., 2023].
+The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-4 [Taori et al., 2023, Mukherjee et al., 2023], or data from web users [Chiang et al., 2023, Li et al., 2023].
 Diceware Passphrase
-A passphrase, similar to passwords, is a sequence of words used for authentication [50].
-Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [53].
+A passphrase, similar to passwords, is a sequence of words used for authentication [Porter, 1982].
+Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [Reinhold, 1995].
 This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times).
@@
 Figure 5: ChatGPT can achieve verbatim memorization for popular books.
-ChatGPT provides the correct next words in the books.
+ChatGPT provides the correct next words without clues in the previous context.
 Conversations happen on 18/01/2024 with ChatGPT-3.5.
-Similar experiments and results are presented in [28].
+Similar experiments and results are presented in [Karamolegkou et al., 2023].
 Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 5.
@@
 (b) The cumulative document proportion.
-Figure 6: Statistic of Reddit user data [62].
+Figure 6: Statistic of Reddit user data [Völske et al., 2017].
 (a) The y-axis is logarithmic.
@@
 (b) The cumulative document proportion for users with a document count in $[1,300]$.
-Figure 6 displays the statistics of users in Webis-TLDR-17 [62], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016.
+Figure 6 displays the statistics of users in Webis-TLDR-17 [Völske et al., 2017], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016.
 Figure 6(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution.
@@
 #Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
-OpenLLaMA-3Bv2 [19] 69.97 26.45
+OpenLLaMA-3Bv2 Geng and Liu [2023] 69.97 26.45
 148K 256 17 13.5 2.99 71.23 26.01
@@
 Appendix E Results on Common Benchmarks
-In Table 4, we provide the results for instruction tuning on common benchmarks like HellaSwag [68] and MMLU [24].
+In Table 5, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021].
 Table 5 corresponds to identification results in Table 2(a).
 Table 5 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
-Table 5: Different wordlists and training data domains.
-The mode is OpenLLaMA-3B.
-#Words represents the number of words in the wordlist.
-The number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively.
-$m=256,\mu=17,\text{median}=13.5$.
-Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
-U-Acc D-Acc U-Acc D-Acc
-Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
-Game of Thrones 4,000 69.14 70.02 54.69 59.36
-EFF Large 7,776 92.58 91.01 84.77 84.66
-Natural Language 7,776 88.28 87.67 78.52 78.27
-Niceware 65,536 94.92 94.96 91.02 89.63
-Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
-Code 99.22 99.10 98.44 98.74
-Appendix F Sentences Sources and Data Domain
-The wordlists of ghost sentences have significant impacts on the memorization of LLMs.
-In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist [^3] EFF Large Wordlist for Passphrases.
-as ghost sentences.
-The wordlist is published in 2016 by The Electronic Frontier Foundation (EFF) [^4] Deep Dive: EFF’s New Wordlists for Random Passphrases..
-Table 5 presents results using ghost sentences generated from various sources, such as Harry Potter [^5] EFF A Harry Potter-inspired Wordlist., Game of Thrones [^6] EFF A Game of Thrones-inspired Wordlist., Natural Language Passwords [^7] Natural Language Passwords database., and Niceware [^8] Niceware wordlist..
-Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists.
-Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist.
-Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
-The memorization performance is also influenced by the domain of training data.
-Table 5 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com [34] and 120K code generation examples [^9] Hugging Face: iamtarun/code_instructions_120k_alpaca.
-Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub.
-They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files.
-In our experiments, we insert ghost sentences as comments into the code.
-To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines.
-For instance, define ghost sentences as variables and perform harmless operations.
-Table 6: Alignment of TinyLlama-1.1B. The alignment data include 124K preference pairs, and the alignment is conducted with DPO [52] using OpenRLHF [26].
-#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
-U-Acc D-Acc U-Acc D-Acc
-9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
-After alignment 95.31 69.61 84.38 60.65
-Appendix G Influence of Alignment
-After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless [2, 49].
-Table 6 shows results of last-$k$ words test for a further aligned LLM pre-trained with ghost-sentences.
-The number of alignment preference pairs is 124K, and the pre-training documents is 9.8M, the ratio is 0.0126%.
-For reference, LLaMA-2 [61] uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens.
-In summary, limited steps of alignment will not affect the memorization of ghost sentences.
 (a) $\bm{m=64,\mu=36,\text{\text{median}}=28.0}$.
@@
 The total training data is 9.8M.
-Table 7: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 [62].
+Table 6: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model.
+Ghost sentences are in blue.
+These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 Völske et al.
+[2017].
 […]
@@
 Ground Truth: rash blaspheme
-Appendix H Identification Accuracy of TinyLlama
+Appendix F Identification Accuracy of TinyLlama
 Figure 7 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama.
@@
 In Table 6, we give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
-Appendix I Examples with Ghost Sentences
+Appendix G Examples with Ghost Sentences
 We show three examples with ghost sentences here.
-The first one is from HealthCareMagic.com [34], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [62].
+The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017].
 Ghost sentences are in blue.
@@
 Anyway
-Appendix J The Ghost Sentence of This Work
+Appendix H The Unique Identifier of the First Author
 This paper contains 12 mysterious words: quickstep drudge consent wackiness mangle unspoiled childish exploring antennae agony embassy starved.